### v1

# Long-Lived Photon Blockade with Weak Optical Nonlinearity

Abstract
In conventional photon blockade, the occupation of a cavity mode by more than one photon is suppressed via strong optical nonlinearity. An alternative, called unconventional photon blockade, can occur under weak nonlinearity by relying on quantum interference between fine-tuned cavities. A serious limitation is the very short antibunching time window, orders of magnitude less than the cavity lifetime. We present a method to achieve photon blockade over a large time window of several cavity lifetimes, even exceeding that of conventional photon blockade, while still requiring only weak nonlinearity. This “long-lived photon blockade” (LLPB) occurs when the single-photon Green’s function exhibits a zero at a large cavity loss rate, which is satisfied by an exemplary configuration of four coupled cavities under weak driving. Our analytical results agree well with wavefunction Monte Carlo simulations. The LLPB phenomenon may aid the development of single-photon sources utilizing materials with weak optical nonlinearities.
Introduction—Photon blockade is a quantum effect whereby one photon in a nonlinear resonator blocks the entry of other photons, giving rise to antibunched photon statistics [1, 2, 3, 4]. Conventionally, photon blockade requires the optical nonlinearity within a cavity to exceed its loss: i.e., $\alpha\gg\gamma$, where $\alpha$ characterizes the strength of the nonlinearity and $\gamma$ is the cavity decay rate. This can be interpreted in terms of a two-photon state being driven off-resonance by the nonlinearity, shifting its energy by more than the cavity linewidth [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Achieving this in the optical regime (including telecom wavelengths) is challenging, as material nonlinearities tend to be very weak relative to cavity linewidths. Systems featuring exceptionally strong optical nonlinearities are required, such as exciton-polaritons [24, 25], trions [26], Rydberg excitons [27], and Moire excitons [28].
The strong-nonlinearity condition may be bypassed via the phenomenon of unconventional photon blockade (UPB), which uses destructive interference between two or more nonlinear cavities (or cavity modes) to cancel the two-photon amplitudes [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. It is commonly believed that UPB requires the inter-cavity coupling rate $J$ to satisfy [29]
$J\gg\gamma.$ (1)
This seems to make intuitive sense: for interference to play a role, the photons must be able to hop many times before leaking away. In this regime, it then turns out that even a weak nonlinearity, $\alpha\ll\gamma$, can produce photon antibunching. However, even if the second-order photon correlation $g^{(2)}(\tau)$ vanishes at $\tau=0$, the suppression only holds over a time window of
$\delta\tau\sim 1/J\ll 1/\gamma,$ (2)
due to the Rabi-like oscillations that $J$ induces in the time-dependent state amplitudes. This is much smaller than the antibunching time window of $\delta\tau\sim 1/\gamma$ for conventional photon blockade, and poses a serious practical obstacle to the realization and exploitation of UPB. For example, a proposal for UPB at telecom wavelengths [37] requires $J\approx 20\gamma$ and yields an antibunching time window of $\sim 100\,\text{ps}$, around the resolution limit for current photodetectors. In such two-cavity models, the UPB condition reduces to $\gamma^{3}\propto\alpha J^{2}$, implying that the cavity quality factors satisfy $Q^{3}\propto\omega_{0}^{3}\delta\tau^{2}/\alpha$, where $\omega_{0}$ is the cavity resonance frequency [31]. Given a lower bound on $\delta\tau$ imposed by detector resolution limits, this condition sets a minimum requirement on $Q$ that is unfavorable in the optical regime, where $\omega_{0}$ is large and optical nonlinearities are typically weak. Flayac and Savona have suggested that with one-way dissipative inter-cavity couplings, the rapid oscillations in $g^{(2)}(\tau)$ can be avoided [51]; however, such couplings are highly challenging to implement experimentally.
In this paper, we show that interference-assisted photon blockade does not require the strong-coupling condition (1), even with weakly nonlinear cavities and ordinary couplings. Using a Green’s function formalism, we analyze the steady-state behavior of coupled nonlinear cavities in the weak-driving regime [31], and show that antibunching solutions appear in the vicinity of 1-photon dark states, where the 1-photon Green’s function vanishes in a specific cavity. In the two-cavity systems where UPB was first discovered [29, 30, 31], as well as more complicated coupled-cavity configurations [50], such dark states occur only at $\gamma=0$, giving rise to UPB solutions with nonzero but small $\gamma$ satisfying Eq. (1). However, certain coupled-cavity configurations host other families of dark states at large $\gamma$ (relative to $J$). We demonstrate this with a four-cavity system, showing that it produces a qualitatively different form of antibunching that we call long-lived photon blockade (LLPB). Unlike the earlier proposal involving one-way dissipative couplings [51], our four-cavity setup involves only standard (i.e., two-way and Hermitian) inter-cavity couplings. The antibunching condition $g^{(2)}(0)=0$ is satisfied alongside weak coupling ($J\ll\gamma$) and weak nonlinearity ($\alpha\ll\gamma$). Moreover, $g^{(2)}(\tau)$ is suppressed over a time window $\delta\tau\approx 8/\gamma$, two orders of magnitude wider than a comparable two-cavity system exhibiting UPB, and even larger than for conventional photon blockade under strong nonlinearity. Thus, it may be possible to achieve photon blockade in the visible regime with weakly nonlinear optical media.
Model—We consider coupled nonlinear cavities, indexed by $i=1,\dots$, under a weak coherent drive. Assuming for simplicity that all the cavities have identical parameters, the cavity Hamiltonian in the frame co-rotating with the driving field, excluding losses, is
$\mathcal{H}_{0}=\sum_{i\neq j}J_{ij}a^{\dagger}_{i}a_{j}+\sum_{i}\Delta a^{\dagger}_{i}a_{i}+\alpha\sum_{i}a^{\dagger}_{i}a^{\dagger}_{i}a_{i}a_{i}.$ (3)
Here, $\hbar=1$, $\Delta$ is the cavity detuning relative to the driving frequency, $\alpha$ is the nonlinearity strength (Kerr coefficient), $J_{ij}$ is the coupling between cavities $i$ and $j$, and $a_{i}$ is the photon annihilation operator for cavity $i$. We assume no magneto-optic activity, so $J_{ij}=J_{ji}\in\mathbb{R}$ with a suitable gauge.
We assume a single-input, single-output setup where a cavity $d$ is driven, and we look at the photon populations in a “readout” cavity (which may or may not be $d$). The driving Hamiltonian is $\mathcal{H}_{d}=F_{d}a_{d}^{\dagger}+F^{*}_{d}a_{d}$, where $F_{d}$ is the complex driving amplitude. (The following derivation can also be straightforwardly generalized to the case of multiple coherently driven cavities.) The density matrix $\rho$ is then governed by the Lindblad master equation [52]
$i\frac{d\rho}{dt}=[\mathcal{H}_{0}+\mathcal{H}_{d},\rho]+\frac{i\gamma}{2}\sum_{i}(2a_{i}\rho a^{\dagger}_{i}-a^{\dagger}_{i}a_{i}\rho-\rho a^{\dagger}_{i}a_{i}),$ (4)
where the terms proportional to the cavity loss rate $\gamma$ describe environmental noise. In the weak drive limit [31], stochastic quantum jumps—the first term in parentheses—can be ignored, and the problem reduces to a dissipative Schrödinger equation,
$i\frac{d|\psi\rangle}{dt}$ $=\big{[}\mathcal{H}(z)+\mathcal{H}_{d}\big{]}|\psi\rangle,$ (5)
$\mathcal{H}(z)$ $\equiv\sum_{i\neq j}J_{ij}a^{\dagger}_{i}a_{j}+\sum_{i}za^{\dagger}_{i}a_{i}+\alpha\sum_{i}a^{\dagger}_{i}a^{\dagger}_{i}a_{i}a_{i},$ (6)
$z$ $\equiv\Delta-i\gamma/2,$ (7)
where the non-Hermitian effective Hamiltonian $\mathcal{H}(z)$ is derived from (3) by adding loss to each site.
We seek a solution that is stationary (i.e., co-rotating with the drive, $d/dt\rightarrow 0$). We truncate to superpositions of only 1- and 2-photon states, and derive the following expression for the equal-time second-order correlation function:
$g_{ij}^{(2)}$ $=\frac{\langle a^{\dagger}_{j}a^{\dagger}_{i}a_{i}a_{j}\rangle}{\langle a^{\dagger}_{j}a_{j}\rangle\langle a^{\dagger}_{i}a_{i}\rangle}=\big{|}f_{ij}(z)\big{|}^{2},$ (8)
$f_{ij}(z)$ $=\frac{\sqrt{2}\langle i,j|\bar{\psi}^{(2)}\rangle}{\langle i|\bar{\psi}^{(1)}\rangle\langle j|\bar{\psi}^{(1)}\rangle},$ (9)
where $|i\rangle=a_{i}^{\dagger}|\text{vac}\rangle$ denotes the state with one photon in cavity $i$ and $|i,j\rangle\equiv|i\rangle\otimes|j\rangle$. This generalizes a result reported by us for a specific lattice model [50]; the details of the derivation are given in the Supplemental Materials [53]. In Eq. (8), the expectation values are taken over the aforementioned stationary state, while in Eq. (9) $|\bar{\psi}^{(1)}\rangle$ and $|\bar{\psi}^{(2)}\rangle$ denote projections of the steady-state wavefunction to the 1- and 2-photon subspaces respectively.
Using Eq. (9), we can construct the following account of how photon antibunching arises. By fine-tuning the system parameters, it may be possible to satisfy $\langle i|\bar{\psi}^{(1)}\rangle=0$, which we call a 1-photon dark state, for some cavity $i$. In the absence of nonlinearity, such a condition also implies a 2-photon dark state $\langle i,j|\bar{\psi}^{(2)}\rangle=0$, since $|\bar{\psi}^{(2)}\rangle=|\bar{\psi}^{(1)}\rangle\otimes|\bar{\psi}^{(1)}\rangle/\sqrt{2}$. We then find that even though the numerator and denominator of (9) both vanish, $f_{ij}(z)$ does not vanish; in fact, $g_{ij}^{(2)}=1$ for all $z$. A weak nonlinearity, however, can induce a small misalignment between the 1-photon and 2-photon dark states. Given a 1-photon dark state at $z_{0}$, the 2-photon dark state condition is shifted to a nearby value of $z_{0}+\delta z$. At this point, the numerator of (9) vanishes while the denominator is nonzero. This is our desired photon antibunching point.
It is also helpful to introduce the 1-photon and non-interacting 2-photon Green’s functions,
$G(z)$ $\equiv\sum_{n}\frac{|\varphi_{n}\rangle\langle\varphi_{n}|}{z+\epsilon_{n}},$ (10)
$G^{(2)}(z)$ $\equiv\sum_{mn}\frac{|\varphi_{m},\varphi_{n}\rangle\langle\varphi_{m},\varphi_{n}|}{2z+\epsilon_{m}+\epsilon_{n}},$ (11)
where $\{|\varphi_{n}\rangle\}$ is the eigenbasis of the coupling matrix, $\left[J_{ij}\right]|\varphi_{n}\rangle=\epsilon_{n}|\varphi_{n}\rangle$, and $|\varphi_{m},\varphi_{n}\rangle\equiv|\varphi_{m}\rangle\otimes|\varphi_{n}\rangle$. Using these and assuming $\alpha$ is perturbative, Eq. (9) can be expressed as
$f_{ij}(z)$ $=1-2\alpha\sum_{k}\frac{G_{kd}^{2}}{G_{id}G_{jd}}\langle i,j|G^{(2)}|k,k\rangle.$ (12)
The time-delayed correlation function $g_{ij}^{(2)}(\tau)$ can be calculated by upgrading the $a_{i}$’s and $a_{i}^{\dagger}$’s to Heisenberg picture operators [53]. For $\tau>0$, we find that the non-interacting 2-photon Green’s function in Eq. (12) is simply replaced by
$G^{(2)}(z,\tau)$ $\equiv e^{-iz\tau}\sum_{mn}\frac{|\varphi_{m},\varphi_{n}\rangle\langle\varphi_{m},\varphi_{n}|}{2z+\epsilon_{m}+\epsilon_{n}}e^{-i\epsilon_{n}\tau}.$ (13)
This analytical form of $g_{ij}^{(2)}(\tau)$ has not previously been reported, to our knowledge. Below, we will use it to revisit UPB, and then demonstrate the LLPB phenomenon.
Conventional blockade—First, let us consider photon blockade in one cavity. In this case, we can bypass the above perturbative procedure and directly find the Green’s functions, obtaining [53]
$g_{11}^{(2)}(\tau)=\frac{1}{1+4\alpha^{2}/\gamma^{2}}+\frac{4\alpha^{2}/\gamma^{2}}{1+4\alpha^{2}/\gamma^{2}}\left(1-e^{-\frac{\gamma}{2}\tau}\right)^{2}.$ (14)
The condition $g_{11}^{(2)}(0)\approx 0$ requires $\alpha\gg\gamma$. The corresponding time window, estimated as the interval over which $g^{(2)}(\tau)<0.5$, is $\delta\tau=-4\rm{ln}(1-\sqrt{0.5})/\gamma\approx 5/\gamma$.
UPB with two cavities—For a two-cavity system, with $J_{12}=J_{21}=J\in\mathbb{R}^{+}$, the coupling matrix has eigenvalues $\epsilon_{1,2}=\pm J$ and the eigenbasis $|\varphi_{1,2}\rangle=2^{-1/2}(|1\rangle\pm|2\rangle)$. Eqs. (10)–(12) predict antibunching in the driven cavity ($d=1$), $g_{11}^{(2)}(0)=0$, if and only if $z^{3}\approx-\alpha J^{2}/2$ [53]. The three roots correspond to three 2-photon dark states, and are depicted by blue circles in Fig. 1(a). They lie close to the 1-photon dark state at $z=0$, for which $\langle 1|\bar{\psi}^{(1)}\rangle\propto G_{11}=z(z^{2}-J^{2})^{-1}=0$, as indicated by the red cross in the figure. We focus on the root with positive cavity loss, for which
$\gamma=\sqrt{3}\Delta\propto J\sqrt{\alpha/\gamma}.$ (15)
Unlike the one-cavity case, this is achievable for $\alpha\ll\gamma$, but only if we enter the strong-coupling regime (1), as previously noted [29, 31]. Moreover, from Eq. (13),
$g_{11}^{(2)}(\tau)$ $\approx\left(1-e^{-\frac{\gamma}{2}\tau}\cos(\Delta\tau)\cos(J\tau)\right)^{2}.$ (16)
The fast oscillation in $\cos(J\tau)$ yields a small time window of $\delta\tau\sim 1/J$, tied to the large coupling eigenvalues ±$\pm J$.
Photon antibunching in this model follows the mechanism laid out by Eqs. (8)–(12). The 2-photon dark state is perturbed away from the 1-photon dark state, and since the latter occurs at $z=0$, the solution we arrive at involves small $\gamma$ (relative to $J$). The lattice model studied in Ref. 50 also works similarly.
Figure 1: (a) Schematic of the zeros (circles) and pole (cross) of $f_{11}(z)$ in the complex $z$ plane, for a two-cavity model. The pole occurs at $z=0$, and UPB occurs at the zeros. (b) Schematic of a four-cavity setup that can exhibit long-lived photon blockade (LLPB). (c) Distribution of zeros (circles) and pole (cross) of $f_{22}(z)$ for the four-cavity model of (b). This illustration is not drawn to scale. (d) Equal-time two-photon correlation function $g^{(2)}(0)$ for the four-cavity system versus cavity detuning $\Delta$ and loss rate $\gamma$. Upper plot: analytical results obtained from Eqs. (20)–(22). Near the pole (central red spot), there are two zeros of $g^{(2)}(0)$ (blue spots) corresponding to LLPB. The model parameters are $k=16$, $J=0.1227$, $J^{\prime}=0.02454$, and $\alpha=0.001227$, chosen so that one of the LLPB points occurs at $\gamma=1$. Lower plot: WFMC simulation results, obtained by directly solving Lindblad master equation with the same parameters and $F_{d}=10^{-5}$.
Long-Lived Photon Blockade—The above scenario can be bypassed if we find a configuration of cavities with a 1-photon dark state at large $\gamma$. Specifically, we want the linear single-photon Green’s function to satisfy $G_{id}(z)=0$ for some $i,d$, with $\gamma=-2\mathrm{Im}(z)\gg\mathrm{max}\,|J_{ij}|$. This turns out to be impossible for $N=2$ and $N=3$ (though if we relax the assumption that the cavities are identical, we can find a zero for $N=3$ if one cavity is completely lossless); for details, see the Supplemental Materials [53]. For $N=4$, however, the condition can be satisfied.
Consider the configuration shown in Fig. 1(b), with four cavities arranged in a ring with couplings $J_{12},J_{23},J_{34},J_{41}\in\mathbb{R}^{+}$; the driven cavity is $d=1$, and photons are extracted from cavity $i=2$. A Dyson series expansion in the weak-coupling regime [54] indicates that $G_{12}(z)=G_{21}(z)$ can have a zero that meets the above conditions if $J_{12}\ll J_{23},J_{34},J_{41}$. Specifically, we consider
$[J_{ij}]=\begin{pmatrix}0&J^{\prime}/k&0&J\\ J^{\prime}/k&0&J&0\\ 0&J&0&J^{\prime}\\ J&0&J^{\prime}&0\end{pmatrix},$ (17)
where $k\gg 1$. Using Eqs. (6) and (10),
$G_{21}(z)$ $=-\frac{J^{\prime}}{k}(z+z_{0})(z-z_{0})/\det\mathcal{H},$ (18)
$z_{0}$ $\equiv-iJ\sqrt{k-(J^{\prime}/J)^{2}}.$ (19)
At $z=z_{0}$, $G_{21}$ has a zero and hence $f_{22}$ has a pole (we ignore $z=-z_{0}$, which involves putting gain in the cavities). From Eqs. (7) and (19), we have $\Delta=0$ and $\gamma\approx 2\sqrt{k}J$, consistent with the weak-coupling regime.
Next, we search near $z_{0}$ for a zero of $f_{22}(z)$. Substituting Eq. (18) back into Eq. (12), we find that for $z\approx z_{0}$,
$f(z)$ $\approx\frac{(z-z_{0}-\delta z)(z-z_{0}+\delta z)}{(z-z_{0})^{2}},$ (20)
$\delta z^{2}$ $=-\dfrac{k^{2}\alpha}{2z_{0}^{2}J^{\prime 2}}\left[{\sum}_{i}G^{(2)}_{22ii}\,\big{(}G_{i1}\,\cdot\,\det\mathcal{H}\big{)}^{2}\right]_{z=z_{0}}$ (21)
where $G^{(2)}_{22ii}\equiv\langle 2,2|G^{(2)}(z)|i,i\rangle$. By further assuming $J\gg J^{\prime}$, it is approximated as [53]
$\delta z\approx\pm{\frac{\sqrt{38}}{16}}\sqrt{\alpha\gamma}e^{-i\pi/4}.$ (22)
LLPB is predicted to occur at $z\approx z_{0}\pm\delta z$, as shown schematically in Fig. 1(c).
In the upper plot of Fig. 1(d), we plot $g_{22}^{(2)}(0)$ versus the cavity parameters $\Delta$ and $\gamma$ using Eqs. (20)–(22). The parameters, stated in the figure caption, satisfy $\alpha\ll\gamma$ (weak nonlinearity) and $J^{\prime}/k\ll J^{\prime}\ll J\ll\gamma$ (weak coupling). For comparison, the lower plot of Fig. 1(d) shows the results obtained based on wavefunction Monte Carlo (WFMC) [55, 56, 57, 58] simulations, which directly solve the Lindblad master equation (4) using $F_{d}=10^{-5}$, with all other parameters kept the same. It is worth noting that the WFMC simulations give $g_{22}^{(2)}(0)\ll 1$, but not exactly zero, in the weak-drive regime. This value of $g_{22}^{(2)}(0)$ increases with $F_{d}$, alongside an increase in the mean photon occupation number in the signal cavity [53].
Figure 2: Time-delayed second-order correlation function $g^{(2)}(\tau)$ for a four-cavity system with LLPB (green curve), a two-cavity system with UPB (blue curve), and a one-cavity system with conventional photon blockade (orange dashes). All results are obtained using WFMC simulations. The model parameters for the four-cavity LLPB setup are the same as in Fig. 1, with $\Delta=0.009571$ and $\gamma=1$. The other two cases are tuned for photon blockade at the same $\gamma$. For the two-cavity system, we choose the same nonlinearity strength $\alpha=0.001227$, for which UPB occurs at $J\approx 17.67$ and $\Delta\approx 0.2915$. For the one-cavity system, we set $\alpha=10$ and $\Delta=0.02491$.
Antibunching time window—In Fig. 2, the green curve shows the time-delayed second-order correlation function $g^{(2)}(\tau)$ for the above four-cavity system, at one of its LLPB points (at $\gamma=1$). The photon antibunching time window is approximately $8/\gamma$. This plot is obtained from WFMC simulations, but the analytical theory gives almost identical results. The analytic approximation for this result can be derived for $J^{\prime}\ll J$ by approximating the eigenvalues and eigenvectors of the coupling Hamiltonian to second order in $J^{\prime}$. The full result is given in Supplemental Materials [53]. If we look into the short time scale $\tau\sim 1/\gamma$ and expand $g^{(2)}_{22}(\tau)$ as polynomial of $\gamma\tau$, assuming negligible $\alpha$, we find
$g_{22}^{(2)}(\tau)$ $\approx(\gamma\tau)^{4}/64.$ (23)
This small $\tau$ expansion is order-4 and fundamentally distinct from the quadratic conventional photon blockade case $g^{(2)}(\tau)\approx(\gamma\tau)^{2}/4$, by Eq. (14).
For comparison, the blue curve in Fig. 2 shows $g^{(2)}(\tau)$ for a comparable two-cavity UPB setup [29]. Here we take $\alpha=0.001227$, the same as in the four-cavity LLPB case. For the same decay rate $\gamma=1$, two-cavity UPB is achieved for $J\approx 17.67$, and the strong coupling results in rapid oscillations of $g^{(2)}(\tau)$. The central antibunching window, estimated as the window over which $g^{(2)}(\tau)<0.5$, is around 90 times smaller than in the LLPB case.
As a further comparison, the orange dashes in Fig. 2 show $g^{(2)}(\tau)$ for a one-cavity setup with $\gamma=1$ and $\alpha=10$. (Note that $g^{(2)}(0)$ does not reach precisely zero under conventional photon blockade; we simply choose $\alpha$ so that it is visibly suppressed.) The $g^{(2)}(\tau)$ curves for both LLPB and conventional photon blockade have no oscillations, unlike the two-cavity UPB case. Remarkably, the LLPB curve has a significantly flatter bottom near $\tau=0$ than the curve for conventional photon blockade, and the time window is larger by a factor of around 1.68. This can be understood from Eq. (23); the expansion leading term is of order-4 in $\gamma\tau$, approximately $(\gamma\tau)^{4}/64$, compared to the quadratic term $\approx(\gamma\tau)^{2}/4$ for the single-cavity case. This characteristic results in a more gradual change in the dip, which may be advantageous for pulsed device operations [37, 39].
Conclusion—We have re-analyzed the origins of unconventional photon blockade (UPB), in which photon antibunching is achieved by quantum interference in weakly nonlinear coupled cavities [29]. A major problem with UPB has hitherto been the rapid oscillation in $g^{(2)}(\tau)$ due to strong inter-cavity coupling [37, 39]. We show that this can be overcome by designing a coupled-cavity system so that the linear single-photon Green’s function exhibits a zero (a 1-photon dark state) at a large cavity loss rate. This leads to a “long-lived photon blockade” (LLPB) scenario whereby $g^{(2)}(\tau)\propto\tau^{4}$ and the time window exceeds that of conventional photon blockade under strong nonlinearity. It is remarkable that shortening the cavity lifetime can effectively lengthen the antibunching time window, which can be regarded as a fresh example of non-Hermiticity (in the form of dissipation) producing unexpected physical behaviors [59, 60]. Our proposal does not need one-way inter-cavity couplings or other exotic requirements [51, 39], and should be highly compatible with quantum optics experiments.
Our theoretical framework offers ample scope for further generalization and optimization. For instance, in the four-cavity configuration of Fig. 1(b), the couplings are described by just three parameters for simplicity, but they can instead be individually optimized to achieve an even larger $g^{(2)}(\tau)$ time window. Coherently driving multiple cavities also offers additional possibilities for achieving LLPB with other cavity configurations. Possible implementation challenges may include the need for fine-tuning and ensuring the stability of the inter-cavity couplings and other system parameters, which is a problem shared with the original UPB setup and all other related schemes to date.
Supplemental Materials
S1 Derivation of delayed second-order correlations under weak driving
As mentioned in the main text, the steady state for a system of driven nonlinear cavities, in the limit where the driving is weak, can be obtained from a dissipative effective Schrödinger equation. Following Ref. [31], under the weak-driving assumption the Hilbert space is divided into subspaces of different photon number, and the wavefunction is written as a superposition of contributions from each subspace:
$|\psi\rangle=|\psi^{(0)}\rangle+|\psi^{(1)}\rangle+|\psi^{(2)}\rangle+\ldots$ (S1)
Under weak driving, the amplitudes in higher photon number subspaces should be negligible. Hence, we solve the Schrödinger equation for each subspace recursively [31], truncating at a given photon number bound:
$i\frac{d}{dt}|\psi^{(k)}\rangle$ $=\mathcal{H}|\psi^{(k)}\rangle+F_{d}a^{\dagger}_{d}|\psi^{(k-1)}\rangle+F_{d}^{*}a_{d}|\psi^{(k+1)}\rangle$
$\approx\mathcal{H}|\psi^{(k)}\rangle+F_{d}a^{\dagger}_{d}|\psi^{(k-1)}\rangle.$ (S2)
The steady-state wavefunction $|\bar{\psi}\rangle$ can be found by setting the time derivative in each subspace to zero. We obtain
$|\bar{\psi}^{(k)}\rangle=-F_{d}\mathcal{H}^{-1}a^{\dagger}_{d}|\bar{\psi}^{(k-1)}\rangle,$ (S3)
with $|\bar{\psi}^{(0)}\rangle=|\psi^{(0)}\rangle$ being the vacuum state.
For the single-photon subspace, the effective Hamiltonian is $\mathcal{H}=[J_{ij}]+z\mathcal{I}$. This subspace is spanned by the eigenvectors of the coupling matrix, i.e.,
$[J_{ij}]|\varphi_{n}\rangle=\epsilon_{n}|\varphi_{n}\rangle.$ (S4)
Supposing that one cavity $d$ is driven, (S3) reduces to
$|\bar{\psi}^{(1)}\rangle=-F_{d}G|d\rangle,\quad\textrm{ where }G\equiv\mathcal{H}^{-1}=\sum_{n}\frac{|\varphi_{n}\rangle\langle\varphi_{n}|}{z+\epsilon_{n}}.$ (S5)
In this work, we truncate the calculation at two photons. For the 2-photon subspace, we use the basis formed by tensor products $|\varphi_{mn}\rangle=|\varphi_{m}\rangle\otimes|\varphi_{n}\rangle$. We then calculate the steady-state amplitude in this subspace using a perturbation expansion to first order in the nonlinearity strength $\alpha$ [50]:
$|\bar{\psi}^{(2)}\rangle$ $\approx|\bar{\psi}_{0}^{(2)}\rangle+\alpha|\bar{\psi}_{1}^{(2)}\rangle$
$=\frac{1}{\sqrt{2}}|\bar{\psi}^{(1)}\rangle\otimes|\bar{\psi}^{(1)}\rangle-\sqrt{2}\alpha F_{d}^{2}\sum_{k}G_{kd}^{2}G^{(2)}|k,k\rangle,$ (S6)
where $G_{kd}\equiv\langle k|G|d\rangle$ is the $k,d$ element of the single-photon Green’s function and
$G^{(2)}\equiv\sum_{mn}\frac{|\varphi_{nm}\rangle\langle\varphi_{mn}|}{2z+\epsilon_{m}+\epsilon_{n}}$ (S7)
is the linear two-photon Green’s function.
The second-order correlation function between sites $i$ and $j$, with time delay $\tau\geq 0$, is defined as
$g^{(2)}_{ij}(\tau)=\lim\limits_{t\rightarrow+\infty}\frac{\langle a^{\dagger}_{j}(t)a^{\dagger}_{i}(t+\tau)a_{i}(t+\tau)a_{j}(t)\rangle}{\langle a^{\dagger}_{j}(t)a_{j}(t)\rangle\langle a^{\dagger}_{i}(t+\tau)a_{i}(t+\tau)\rangle},$ (S8)
where the expectation values are taken over the vacuum state. In the Schrödinger representation, this can be transformed to [61]
$g^{(2)}_{ij}(\tau)$ $=\frac{\mathrm{tr}[\bar{\rho}a^{\dagger}_{j}U^{\dagger}(\tau)a^{\dagger}_{i}a_{i}U(\tau)a_{j}]}{\bar{n}_{i}\bar{n}_{j}}$
$=\frac{\mathrm{tr}[U(\tau)\rho_{j}(0)U^{\dagger}(\tau)a^{\dagger}_{i}a_{i}]}{\bar{n}_{i}\bar{n}_{j}},$ (S9)
where $U(\tau)=\mathrm{exp}(-i\mathcal{H}_{tot}\tau)$ is the time-evolution operator, $\bar{\rho}$ is the steady state density operator, and $\rho_{j}(0)\equiv a_{j}\bar{\rho}a^{\dagger}_{j}$. In the denominator, $\bar{n}_{i}\approx|\langle i|\bar{\psi}^{(1)}\rangle|^{2}=|F_{d}G_{id}|^{2}$ under the weak drive approximation. The numerator can be calculated as $\mathrm{tr}[\rho_{j}(\tau)a^{\dagger}_{i}a_{i}]$, where
$\begin{cases}\rho_{j}(0)=a_{j}\bar{\rho}a^{\dagger}_{j}\\ \\ \rho_{j}(\tau)=U(\tau)\rho_{j}(0)U^{\dagger}(\tau).\end{cases}$ (S10)
This is equivalent to calculating the $i$-th site’s photon population at time $\tau$, using the modified initial state $\rho_{j}(0)$.
Now we convert the density operator back into the form $\rho_{j}(\tau)\equiv|\psi_{j}(\tau)\rangle\langle\psi_{j}(\tau)|$, and plug it into the numerator in Eq. (S9):
$\mathrm{tr}[U(\tau)\rho_{j}(0)U^{\dagger}(\tau)a^{\dagger}_{i}a_{i}]=\mathrm{tr}[\rho_{j}(\tau)a^{\dagger}_{i}a_{i}]\approx|\langle i|\psi_{j}^{(1)}(\tau)\rangle|^{2}.$ (S11)
The Schrödinger equation corresponding to Eq. (S10) is
$\begin{cases}|\psi_{j}(0)\rangle=a_{j}|\bar{\psi}\rangle\\ \\ \displaystyle i\frac{d}{d\tau}|\psi_{j}(\tau)\rangle=\mathcal{H}_{tot}|\psi_{j}(\tau)\rangle.\end{cases}$ (S12)
We then recursively solve Eq. (S12) via Eq. (S2):
$|\psi_{j}^{(k)}(0)\rangle$ $=a_{j}|\bar{\psi}^{(k+1)}\rangle,$
$i\frac{d}{d\tau}|\psi_{j}^{(k)}(\tau)\rangle$ $=\mathcal{H}|\psi_{j}^{(k)}(\tau)\rangle+\mathcal{H}_{+}|\psi_{j}^{(k-1)}(\tau)\rangle.$ (S13)
For the single-photon ($k=1$) subspace,
$|\psi_{j}^{(1)}(0)\rangle$ $=a_{j}|\bar{\psi}^{(2)}\rangle,$
$i\frac{d}{d\tau}|\psi_{j}^{(1)}(\tau)\rangle$ $=\mathcal{H}|\psi_{j}^{(1)}(\tau)\rangle+F_{d}a^{\dagger}_{d}a_{j}|\bar{\psi}^{(1)}\rangle$
$=\mathcal{H}|\psi_{j}^{(1)}(\tau)\rangle+F_{d}\langle j|\bar{\psi}^{(1)}\rangle|d\rangle,$ (S14)
where we make use of the fact that $|\psi_{j}^{(0)}(\tau)\rangle=|\psi_{j}^{(0)}(0)\rangle=a_{j}|\bar{\psi}^{(1)}\rangle$ is the vacuum state. The solution to Eq. (S14) is
$|\psi_{j}^{(1)}(\tau)\rangle=|\psi_{j}^{(1)}(+\infty)\rangle+e^{-i\mathcal{H}\tau}[|\psi_{j}^{(1)}(0)\rangle-|\psi_{j}^{(1)}(+\infty)\rangle].$ (S15)
The steady state $|\psi_{j}^{(1)}(+\infty)\rangle$ can be obtained by setting the time derivative to zero in Eq. (S14):
$0=$ $\mathcal{H}|\psi_{j}^{(1)}(+\infty)\rangle+F_{d}\langle j|\bar{\psi}^{(1)}\rangle|d\rangle,$
$|\psi_{j}^{(1)}(+\infty)\rangle$ $=-\mathcal{H}^{-1}F_{d}\langle j|\bar{\psi}^{(1)}\rangle|d\rangle=\langle j|\bar{\psi}^{(1)}\rangle|\bar{\psi}^{(1)}\rangle.$ (S16)
The initial state $|\psi_{j}^{(1)}(0)\rangle=a_{j}|\bar{\psi}^{(2)}\rangle$ is directly obtained by referring to Eq. (S6):
$|\psi_{j}^{(1)}(0)\rangle=a_{j}|\bar{\psi}^{(2)}\rangle$ $=\langle j|\bar{\psi}^{(1)}\rangle|\bar{\psi}^{(1)}\rangle-\sqrt{2}\alpha F_{d}^{2}\sum_{mnk}G_{kd}^{2}\frac{\langle\varphi_{mn}|k,k\rangle}{2z+\epsilon_{m}+\epsilon_{n}}\frac{\langle j|\varphi_{m}\rangle|\varphi_{n}\rangle+\langle j|\varphi_{n}\rangle|\varphi_{m}\rangle}{\sqrt{2}},$
$=\langle j|\bar{\psi}^{(1)}\rangle|\bar{\psi}^{(1)}\rangle-2\alpha F_{d}^{2}\sum_{mnk}\frac{G_{kd}^{2}\langle\varphi_{mn}|k,k\rangle\langle j|\varphi_{m}\rangle}{2z+\epsilon_{m}+\epsilon_{n}}|\varphi_{n}\rangle.$ (S17)
Finally, plugging Eq. (S17) and Eq. (S16) back into Eq. (S15) gives
$\langle i|\psi_{j}^{(1)}(\tau)\rangle=\langle j|\bar{\psi}^{(1)}\rangle\langle i|\bar{\psi}^{(1)}\rangle-2\alpha F_{d}^{2}\sum_{k}G_{kd}^{2}\langle i,j|G^{(2)}(\tau)|k,k\rangle,\$ (S18)
where
$G^{(2)}(\tau)\equiv\sum_{mn}\frac{|\varphi_{mn}\rangle\langle\varphi_{mn}|}{2z+\epsilon_{m}+\epsilon_{n}}e^{-i(z+\epsilon_{n})\tau}.$ (S19)
We finally arrive at a perturbative solution for the delayed second-order correlation function:
$g_{ij}^{(2)}(\tau)$ $=\left|\frac{\langle i|\psi_{j}^{(1)}(\tau)\rangle^{2}}{\bar{n}_{i}\bar{n}_{j}}\right|=\left|\frac{\langle i|\psi_{j}^{(1)}(\tau)\rangle}{\langle i|\bar{\psi}^{(1)}\rangle\langle j|\bar{\psi}^{(1)}\rangle}\right|^{2}$ (S20)
$=\left|1-2\alpha\sum_{k}\frac{G_{kd}^{2}}{G_{id}G_{jd}}\langle i,j|G^{(2)}(\tau)|k,k\rangle\right|^{2}.$ (S21)
S2 Conventional photon blockade
To model conventional photon blockade model, we cannot use the perturbative procedure from the previous section, as $\alpha$ is large. However, Eqs. (S9)–(S15) remain valid, and can be solved directly for a single-cavity system. In this case, the non-driven Hamiltonian is
$\mathcal{H}=za^{\dagger}_{1}a_{1}+\alpha a^{\dagger}_{1}a^{\dagger}_{1}a_{1}a_{1},$ (S22)
and the 1-photon and 2-photon Green’s functions are
$G=\frac{|1\rangle\langle 1|}{z},\quad G^{(2)}=\frac{|1,1\rangle\langle 1,1|}{2z+2\alpha}.$ (S23)
According to Eq. (S3), the steady state wavefunctions are
$|\bar{\psi}^{(1)}\rangle=-\frac{F}{z}|1\rangle,\quad|\bar{\psi}^{(2)}\rangle=\frac{F^{2}}{\sqrt{2}z(z+\alpha)}|1,1\rangle.$ (S24)
Plugging into Eqs. (S14) and Eq. (S15), we obtain
$|\psi_{1}^{(1)}(0)\rangle$ $=a_{1}|\bar{\psi}^{(2)}\rangle=\frac{F^{2}}{z(z+\alpha)}|1\rangle$ (S25)
$|\psi_{1}^{(1)}(+\infty)\rangle$ $=\frac{F^{2}}{z^{2}}|1\rangle$ (S26)
$\Rightarrow\qquad|\psi_{1}^{(1)}(\tau)\rangle$ $=\frac{F^{2}}{z^{2}}\left(1-\frac{\alpha}{\alpha+z}e^{-iz\tau}\right)|1\rangle.$ (S27)
Thus, the second-order correlation function is
$g_{11}^{(2)}(\tau)=\left|\frac{\langle 1|\psi_{1}^{(1)}(\tau)\rangle}{\langle 1|\bar{\psi}^{(1)}\rangle^{2}}\right|^{2}=\left|1-\frac{\alpha}{\alpha+z}e^{-iz\tau}\right|^{2}.$ (S28)
Optimal parameters for photon blockade can be found by minimizing the equal-time second-order correlation:
$g_{11}^{(2)}(0)$ $=\left|\frac{z}{\alpha+z}\right|^{2}=\frac{\Delta^{2}+\frac{\gamma^{2}}{4}}{(\Delta+\alpha)^{2}+\frac{\gamma^{2}}{4}},$ (S29)
$\Delta_{min}$ $=-\frac{1}{2}(\alpha-\sqrt{\alpha^{2}+\gamma^{2}}).$ (S30)
For large nonlinearity $\alpha\gg\gamma$, we can achieve $\Delta_{min}\approx 0$, and the corresponding time-delayed correlation function is
$g_{11}^{(2)}(\tau)$ $=1-\frac{4\alpha^{2}(2e^{-\frac{\gamma}{2}\tau}-e^{-\gamma\tau})}{4\alpha^{2}+\gamma^{2}}$ (S31)
$=\frac{\gamma^{2}}{4\alpha^{2}+\gamma^{2}}+\frac{4\alpha^{2}}{4\alpha^{2}+\gamma^{2}}\left(1-e^{-\frac{\gamma}{2}\tau}\right)^{2}$ (S32)
$\approx\left(1-e^{-\frac{\gamma}{2}\tau}\right)^{2}.$ (S33)
Fig. S1 shows an exemplary plot of $g_{11}^{(2)}(\tau)$ versus $\tau$, comparing Eq. (S32) (solid blue curve) to wavefunction Monte Carlo (WFMC) simulations (black dots). The theoretical results are clearly very accurate.
Figure S1: Time-delayed correlation function $g_{11}^{(2)}(\tau)$ for conventional photon blockade in a single cavity, using the analytical result (S32) (solid blue line) and WFMC simulations (black dots). The system parameters are the same as in the conventional blockade system from Fig. 2 of the main text.
S3 Unconventional photon blockade in two cavities
In this section, we re-derive unconventional photon blockade (UPB) in a two-cavity system using the framework of Section S1. Here, the coupling Hamiltonian is
$[J_{ij}]=\begin{pmatrix}0&J\\ J&0\end{pmatrix},$ (S34)
and its eigenvectors and eigenvalues are
$\begin{pmatrix}|\varphi_{1}\rangle&|\varphi_{2}\rangle\end{pmatrix}=\frac{1}{\sqrt{2}}\begin{pmatrix}-1&1\\ 1&1\end{pmatrix},\quad\begin{pmatrix}\epsilon_{1}\\ \epsilon_{2}\end{pmatrix}=\begin{pmatrix}-J\\ J\end{pmatrix}.$ (S35)
As in the original UPB model [29, 30, 31], we drive and measure at cavity 1. From Eq. (S21),
$g_{ij}^{(2)}(\tau)$ $=\left|1-2\alpha e^{-iz\tau}\sum_{n}A_{ij}(n)e^{-i\epsilon_{n}\tau}\right|^{2},$ (S36)
$A_{ij}(n)$ $=\sum_{km}\frac{G^{2}_{kd}}{G_{id}G_{jd}}\frac{\langle i,j|\varphi_{mn}\rangle\langle\varphi_{mn}|k,k\rangle}{2z+\epsilon_{m}+\epsilon_{n}}.$ (S37)
By inserting Eq. (S35) into Eq. (S37) (note that $G=\sum_{n}(z+\epsilon_{n})^{-1}|\varphi_{n}\rangle\langle\varphi_{n}|$), we have
$A_{11}(1)$ $=\frac{2z^{3}-Jz^{2}+J^{3}}{8z^{3}(z-J)},\quad A_{11}(2)=-\frac{2z^{3}+Jz^{2}-J^{3}}{8z^{3}(z+J)}.$ (S38)
Hence the interfered oscillation terms sum up to
$\sum_{n}A_{11}(n)e^{-i\epsilon_{n}\tau}=\frac{2z^{4}-J^{2}z^{2}+J^{4}}{4z^{3}(z^{2}-J^{2})}\cos J\tau+i\frac{Jz(z^{2}+J^{2})}{4z^{3}(z^{2}-J^{2})}\sin J\tau.$ (S39)
The optimal antibunching condition is
$g_{11}^{(2)}(0)=\left|1-2\alpha\cdot\frac{2z^{4}-J^{2}z^{2}+J^{4}}{4z^{3}(z^{2}-J^{2})}\right|^{2}=0.$ (S40)
This can be satisfied for weak nonlinearity (i.e., small $\alpha$) by taking $J$ to be large, so that
$z=\Delta-\frac{i}{2}\gamma\approx\left(\frac{\alpha J^{2}}{2}\right)^{\frac{1}{3}}e^{-\frac{i\pi}{3}}.$ (S41)
This is consistent with the results of earlier studies, such as Ref. [31]. The corresponding delayed second-order correlation function is
$g_{11}^{(2)}(\tau)$ $=\left|1-e^{-iz\tau}\cos J\tau+iO(\alpha^{\frac{1}{3}})\right|^{2}$
$\approx\left|1-e^{-\frac{\gamma}{2}\tau}e^{-i\Delta\tau}\cos J\tau\right|^{2}$
$\approx\left(1-e^{-\frac{\gamma}{2}\tau}\cos{\Delta\tau}\cos J\tau\right)^{2}.$ (S42)
In the last line, we have used the fact that $\Delta\ll J$ since $(\Delta/J)^{3}\approx(1/16)\alpha/J\ll 1$, and the imaginary part of the slow-varying $e^{-i\Delta\tau}$ can be omitted. In Fig. S2, we compare this analytical formula to WFMC simulation results, using the same UPB settings as in Fig. 2 of the main text.
Figure S2: Time-delayed correlation function $g_{11}^{(2)}(\tau)$ for UPB in a two-cavity system, showing analytical results (solid blue curve) and WFMC simulation results (black dots). The system parameters are the same as in Fig. 2 of the main text.
S4 Long-lived photon blockade
As discussed in the main text, our desired route to photon blockade is to find a situation where $G_{ij}=0$ at a large value of the cavity decay rate (relative to the coupling strength). Let us consider a more general scenario than in the main text, whereby all the couplings $J_{ij}$ and the complex detunings $z_{i}$ are independently tunable. For the two-cavity system, we have
$\mathcal{H}=\begin{pmatrix}z_{1}&J\\ J&z_{2}\end{pmatrix},\hskip 10.0ptG=\frac{1}{z_{1}z_{2}-J^{2}}\begin{pmatrix}z_{2}&-J\\ -J&z_{1}\end{pmatrix}.$ (S43)
Clearly, if $J$ is non-vanishing, a zero of $G$ is only achievable for $z_{1}=0$ or $z_{2}=0$, which does not meet our requirements.
For three cavities,
$\mathcal{H}=\begin{pmatrix}z_{1}&J_{1}&J_{3}\\ J_{1}&z_{2}&J_{2}\\ J_{3}&J_{2}&z_{3}\end{pmatrix},\hskip 10.0ptG=\frac{1}{z_{1}z_{2}z_{3}-z_{1}J_{2}^{2}-z_{2}J_{3}^{2}-z_{3}J_{1}^{2}+2J_{1}J_{2}J_{3}}\begin{pmatrix}z_{2}z_{3}-J_{2}^{2}&J_{2}J_{3}-z_{3}J_{1}&J_{1}J_{2}-z_{2}J_{3}\\ J_{2}J_{3}-z_{3}J_{1}&z_{3}z_{1}-J_{3}^{2}&J_{3}J_{1}-z_{1}J_{2}\\ J_{1}J_{2}-z_{2}J_{3}&J_{3}J_{1}-z_{1}J_{2}&z_{1}z_{2}-J_{1}^{2}\end{pmatrix}.$ (S44)
There are two cases to consider. (i) To create a zero on the diagonal, two complex detunings must be conjugates of each other (e.g., $z_{2}=z_{3}^{*}$ and $|z_{2}|=|z_{3}|=J_{2}$). This requires either that one cavity has gain precisely balanced with another cavity’s loss, or that both cavities are lossless. (ii) To create a zero off-diagonal entry, one complex detuning must be real (e.g., $z_{3}=J_{2}J_{3}/J_{1}$), implying that the cavity must be completely lossless. Either requirement is somewhat impractical (though not impossible) compared to the four-cavity setup.
Let us now analyze the four-cavity system, shown schematically in Fig. 1(b) of the main text.
First, we will develop some intuition about how photon antibunching arises in this system. In the Dyson series expansion for $G_{id}(z)$, each term in the series can be interpreted as a photon trajectory starting at the driven cavity $d$ and ending at the output cavity $i$ [54]. Along a trajectory, each cavity contributes $1/z$ to the Green’s function. For $G_{21}$, the Dyson series up to order $1/z^{4}$ has five terms, corresponding to the following trajectories:
Trajectory Dyson series term
$1\mapsto 2$ $\frac{1}{z}\cdot J_{12}\cdot\frac{1}{z}$
$1\mapsto 2\mapsto 1\mapsto 2$ $\frac{1}{z}\cdot J_{12}\cdot\frac{1}{z}\cdot J_{12}\cdot\frac{1}{z}\cdot J_{12}\cdot\frac{1}{z}$
$1\mapsto 2\mapsto 3\mapsto 2$ $\frac{1}{z}\cdot J_{12}\cdot\frac{1}{z}\cdot J_{23}\cdot\frac{1}{z}\cdot J_{23}\cdot\frac{1}{z}$
$1\mapsto 4\mapsto 1\mapsto 2$ $\frac{1}{z}\cdot J_{41}\cdot\frac{1}{z}\cdot J_{41}\cdot\frac{1}{z}\cdot J_{12}\cdot\frac{1}{z}$
$1\mapsto 4\mapsto 3\mapsto 2$ $\frac{1}{z}\cdot J_{41}\cdot\frac{1}{z}\cdot J_{34}\cdot\frac{1}{z}\cdot J_{23}\cdot\frac{1}{z}$
If we take the sum of these terms and equating it zero, we arrive at the following solution:
$z=\pm i\sqrt{J_{41}^{2}+J_{12}^{2}+J_{23}^{2}+\frac{J_{41}J_{23}J_{34}}{J_{12}}}.$ (S45)
When $J_{12}\ll J_{41},J_{23},J_{34}$, this solution is self-consistent as $\gamma=-2\mathrm{Im}(z)\approx 2\sqrt{J_{41}J_{23}J_{34}/J_{12}}$ is indeed larger than all the coupling terms.
Having established the feasibility of LLPB in this system, let us consider the simplified case $J^{\prime}/J\ll 1$, which locates the pole at $z_{0}=-iJ\sqrt{k-J^{\prime 2}/J^{2}}\approx-iJ\sqrt{k}$ and $\gamma\approx 2\sqrt{k}J$. Then the eigenvectors and eigenvalues of the coupling Hamiltonian $[J_{ij}]$ can be analytically written as second-order perturbation of $J^{\prime}$:
$(|\varphi_{1}\rangle~{}|\varphi_{2}\rangle~{}|\varphi_{3}\rangle~{}|\varphi_{4}\rangle)$ $=\frac{1}{2}\begin{pmatrix}-1&-1&1&1\\ 1&-1&-1&1\\ -1&1&-1&1\\ 1&1&1&1\end{pmatrix}+\frac{J^{\prime}}{8J}\begin{pmatrix}1-\frac{1}{k}+\frac{J^{\prime}}{8J}&-1+\frac{1}{k}+\frac{J^{\prime}}{8J}&1-\frac{1}{k}-\frac{J^{\prime}}{8J}&-1+\frac{1}{k}-\frac{J^{\prime}}{8J}\\ -1+\frac{1}{k}-\frac{J^{\prime}}{8J}&-1+\frac{1}{k}+\frac{J^{\prime}}{8J}&-1+\frac{1}{k}+\frac{J^{\prime}}{8J}&-1+\frac{1}{k}-\frac{J^{\prime}}{8J}\\ -1+\frac{1}{k}+\frac{J^{\prime}}{8J}&-1+\frac{1}{k}-\frac{J^{\prime}}{8J}&1-\frac{1}{k}+\frac{J^{\prime}}{8J}&1-\frac{1}{k}-\frac{J^{\prime}}{8J}\\ 1-\frac{1}{k}-\frac{J^{\prime}}{8J}&-1+\frac{1}{k}-\frac{J^{\prime}}{8J}&-1+\frac{1}{k}-\frac{J^{\prime}}{8J}&1-\frac{1}{k}-\frac{J^{\prime}}{8J}\end{pmatrix},$ (S46)
$\begin{pmatrix}\epsilon_{1}\\ \epsilon_{2}\\ \epsilon_{3}\\ \epsilon_{4}\end{pmatrix}$ $=J\begin{pmatrix}-1\\ -1\\ 1\\ 1\end{pmatrix}+\frac{J^{\prime}}{2}\begin{pmatrix}-1-\frac{1}{k}-\frac{J^{\prime}}{4J}\\ 1+\frac{1}{k}-\frac{J^{\prime}}{4J}\\ -1-\frac{1}{k}+\frac{J^{\prime}}{4J}\\ 1+\frac{1}{k}+\frac{J^{\prime}}{4J}\end{pmatrix}.$ (S47)
The single photon Green’s function to the second order of $J^{\prime}/J$ are given by
$G=\left(\begin{array}[]{cccc}\frac{J^{2}J^{\prime 2}z}{\left(z^{2}-J^{2}\right)^{3}}+\frac{z}{z^{2}-J^{2}}&-\frac{J^{\prime}z^{2}}{k\left(J^{2}-z^{2}\right)^{2}}-\frac{J^{2}J^{\prime}}{\left(J^{2}-z^{2}\right)^{2}}&\frac{JJ^{\prime}z}{k\left(J^{2}-z^{2}\right)^{2}}+\frac{JJ^{\prime}z}{\left(J^{2}-z^{2}\right)^{2}}&\frac{JJ^{\prime 2}z^{2}}{\left(J^{2}-z^{2}\right)^{3}}+\frac{J}{J^{2}-z^{2}}\\ -\frac{J^{\prime}z^{2}}{k\left(J^{2}-z^{2}\right)^{2}}-\frac{J^{2}J^{\prime}}{\left(J^{2}-z^{2}\right)^{2}}&\frac{J^{2}J^{\prime 2}z}{\left(z^{2}-J^{2}\right)^{3}}+\frac{z}{z^{2}-J^{2}}&\frac{JJ^{\prime 2}z^{2}}{\left(J^{2}-z^{2}\right)^{3}}+\frac{J}{J^{2}-z^{2}}&\frac{JJ^{\prime}z}{k\left(J^{2}-z^{2}\right)^{2}}+\frac{JJ^{\prime}z}{\left(J^{2}-z^{2}\right)^{2}}\\ \frac{JJ^{\prime}z}{k\left(J^{2}-z^{2}\right)^{2}}+\frac{JJ^{\prime}z}{\left(J^{2}-z^{2}\right)^{2}}&\frac{JJ^{\prime 2}z^{2}}{\left(J^{2}-z^{2}\right)^{3}}+\frac{J}{J^{2}-z^{2}}&\frac{J^{\prime 2}z^{3}}{\left(z^{2}-J^{2}\right)^{3}}+\frac{z}{z^{2}-J^{2}}&-\frac{J^{2}J^{\prime}}{k\left(J^{2}-z^{2}\right)^{2}}-\frac{J^{\prime}z^{2}}{\left(J^{2}-z^{2}\right)^{2}}\\ \frac{JJ^{\prime 2}z^{2}}{\left(J^{2}-z^{2}\right)^{3}}+\frac{J}{J^{2}-z^{2}}&\frac{JJ^{\prime}z}{k\left(J^{2}-z^{2}\right)^{2}}+\frac{JJ^{\prime}z}{\left(J^{2}-z^{2}\right)^{2}}&-\frac{J^{2}J^{\prime}}{k\left(J^{2}-z^{2}\right)^{2}}-\frac{J^{\prime}z^{2}}{\left(J^{2}-z^{2}\right)^{2}}&\frac{J^{\prime 2}z^{3}}{\left(z^{2}-J^{2}\right)^{3}}+\frac{z}{z^{2}-J^{2}}\\ \end{array}.\right)$ (S52)
Let $G_{12}=0$, we get the pole $z_{0}=-i\sqrt{k}J$. Expanding $z$ near the pole as $z=z_{0}+\delta z$, plugging the eigenvectors, eigenvalues and Green’s function into Eq. (S37), we get:
$2\alpha\sum_{n}A_{22}(n)e^{-i\epsilon_{n}\tau}\approx$ $\frac{i\alpha J}{1024\delta z^{2}}\left\{\cos(J\tau)\left[16J\tau\left(-3k^{2}-10k\right)-304\sqrt{k}+3\frac{J^{\prime 2}\tau}{J}k^{3}\right]\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\left.\sin(J\tau)\left[48k^{2}-144k-3\frac{J^{\prime 2}\tau}{J}k^{5/2}-3\frac{J^{\prime 2}}{J^{2}}k^{3}\right]\right\}.$ (S53)
Requiring $g^{(2)}_{22}(0)=0$ gives
$-\frac{19i\alpha\sqrt{k}J}{64\delta z^{2}}=1\Rightarrow\delta z=\pm\frac{\sqrt{19}}{8}k^{1/4}\sqrt{\alpha J}e^{-\pi i/4}.$ (S54)
Hence, Eq.(S53) becomes
$2\alpha\sum_{n}A_{22}(n)e^{-i\epsilon_{n}\tau}\approx\frac{\sqrt{k}}{304}$$\left\{\cos(J\tau)\left[16J\tau(3k+10)+\frac{304}{\sqrt{k}}-3\frac{J^{\prime 2}\tau}{J}k^{2}\right]\right.$
$+$ $\left.3\sin(J\tau)\left[48-16k+\frac{J^{\prime 2}\tau}{J}k^{3/2}+\frac{J^{\prime 2}}{J^{2}}k^{2}\right]\right\}.$ (S55)
The comparison between Eq. (S55) and WFMC results are shown in Fig. S3.
In the short timescales $J\tau\ll 1$,
$2\alpha\sum_{n}A_{22}(n)e^{-i\epsilon_{n}\tau}\approx 1+\sqrt{k}J\tau+\frac{-152+3J^{\prime 2}k^{2}/J^{2}}{304}J^{2}\tau^{2}$ (S56)
The correlation function $g^{(2)}_{22}(\tau)$ is then given by Eq. (S36) with $z$ given by $z=z_{0}+\delta z$. In the short timescales $J\tau\ll 1$, $g^{(2)}_{22}(\tau)$ behaves as
$g^{(2)}_{22}(\tau)$ $=\left|1-2\alpha e^{-iz\tau}\sum_{n}A_{22}(n)e^{-i\epsilon_{n}\tau}\right|^{2}$
$\approx\left|\frac{(1+i)\sqrt{19}}{8\sqrt{2}}\sqrt{\frac{\alpha}{J}}k^{1/4}J\tau+\frac{1}{128}\left(64k-19i\frac{\alpha}{J}\sqrt{k}\right)J^{2}\tau^{2}\right|^{2}.$ (S57)
[…]
$g^{(2)}_{22}(\tau)\approx\left|\frac{1}{2}kJ^{2}\tau^{2}\right|^{2}\approx(\gamma\tau)^{4}/64.$ (S58)
Figure S3: Time-delayed correlation function $g_{22}^{(2)}(\tau)$ for weak-coupling photon blockade in a four-cavity system, using the full analytical formula (S55) (solid blue curve), and WFMC simulations (black dash). The system parameters are the same as in Fig. 2 of the main text.
S5 Equal-time second order correlation vs. photon occupation number
This section looks into the dependence of the equal-time second order correlation on photon occupation in the signal cavity. Due to the weak coupling $J^{\prime}/k$ between the driven cavity and signal cavity, the photon number in the driven cavity is several orders of magnitude larger than in the signal cavity, making the WFMC simulation with high signal cavity photon occupation difficult.
To overcome this problem and demonstrate the $g^{(2)}(\tau)$ dependence on photon occupation $n$, we use another set of parameters with $k=4$ instead of $k=16$, and $J^{\prime}=J=0.5386$ instead of $J^{\prime}=0.2J$. The other parameters are $\gamma=1$, $\alpha=0.0194$ and $\Delta=0.0652$. Fig. S4(a) shows the resulting comparison between the LLPB system, the two-cavity UPB model (with the same $\gamma$ and $\alpha$), and a conventional photon blockade system (with the same $\gamma$ and large $\alpha=10$). The enlarged antibunching time window and the flat bottom near $\tau=0$ can still be observed. Fig. S4(b) then shows the equal-time second order correlation versus the photon occupation number in the signal cavity.
Figure S4: WFMC simulations of the weakly coupled UPB model for another parameter setup with $k=4$, $\gamma=1$, $J^{\prime}=J=0.5386$, $\alpha=0.0194$ and $\Delta=0.0652$. (a) Comparison with conventional photon blockade and original UPB models. (b) Signal cavity photon occupation number versus equal-time second order correlation.

### v2

@@
 Conventionally, photon blockade requires the optical nonlinearity within a cavity to exceed its loss: i.e., $\alpha\gg\gamma$, where $\alpha$ characterizes the strength of the nonlinearity and $\gamma$ is the cavity decay rate.
-This can be interpreted in terms of a two-photon state being driven off-resonance by the nonlinearity, shifting its energy by more than the cavity linewidth [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
+This can be interpreted in terms of a two-photon state being driven off-resonance by the nonlinearity, shifting its energy by more than the cavity linewidth [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24].
 Achieving this in the optical regime (including telecom wavelengths) is challenging, as material nonlinearities tend to be very weak relative to cavity linewidths.
 Systems featuring exceptionally strong optical nonlinearities are required, such as exciton-polaritons [25, 26], trions [27], Rydberg excitons [28], and Moire excitons [29].
-The strong-nonlinearity condition may be bypassed via the phenomenon of unconventional photon blockade (UPB), which uses destructive interference between two or more nonlinear cavities (or cavity modes) to cancel the two-photon amplitudes [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50].
-It is commonly believed that UPB requires the inter-cavity coupling rate $J$ to satisfy [29]
-$J\gg\gamma.$ (1)
-This seems to make intuitive sense: for interference to play a role, the photons must be able to hop many times before leaking away.
+The strong-nonlinearity condition may be bypassed via the unconventional photon blockade (UPB) phenomenon, which uses destructive interference between two or more nonlinear cavities (or cavity modes) to cancel the two-photon amplitudes [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52].
+It is commonly believed that UPB requires the inter-cavity coupling rate $J$ to satisfy
+$J\gg\gamma,$ (1)
+which seems to make intuitive sense: for interference to play a role, the photons must be able to hop many times before leaking away [30].
 In this regime, it then turns out that even a weak nonlinearity, $\alpha\ll\gamma$, can produce photon antibunching.
@@
 The driving Hamiltonian is $\mathcal{H}_{d}=F_{d}a_{d}^{\dagger}+F^{*}_{d}a_{d}$, where $F_{d}$ is the complex driving amplitude.
-(The following derivation can also be straightforwardly generalized to the case of multiple coherently driven cavities.) The density matrix $\rho$ is then governed by the Lindblad master equation [52]
+(The following derivation can also be straightforwardly generalized to the case of multiple coherently driven cavities.) The density matrix $\rho$ is governed by the Lindblad master equation [54]
 $i\frac{d\rho}{dt}=[\mathcal{H}_{0}+\mathcal{H}_{d},\rho]+\frac{i\gamma}{2}\sum_{i}(2a_{i}\rho a^{\dagger}_{i}-a^{\dagger}_{i}a_{i}\rho-\rho a^{\dagger}_{i}a_{i}),$ (4)
@@
 We seek a solution that is stationary (i.e., co-rotating with the drive, $d/dt\rightarrow 0$).
-We truncate to superpositions of only 1- and 2-photon states, and derive the following expression for the equal-time second-order correlation function:
+Truncating to superpositions of only 1- and 2-photon states yields the following equal-time second-order correlation function:
 $g_{ij}^{(2)}$ $=\frac{\langle a^{\dagger}_{j}a^{\dagger}_{i}a_{i}a_{j}\rangle}{\langle a^{\dagger}_{j}a_{j}\rangle\langle a^{\dagger}_{i}a_{i}\rangle}=\big{|}f_{ij}(z)\big{|}^{2},$ (8)
@@
 where $|i\rangle=a_{i}^{\dagger}|\text{vac}\rangle$ denotes the state with one photon in cavity $i$ and $|i,j\rangle\equiv|i\rangle\otimes|j\rangle$.
-This generalizes a result reported by us for a specific lattice model [50]; the details of the derivation are given in the Supplemental Materials [53].
+The full derivation, which generalizes a previous result based on a specific lattice model [52], is given in the Supplemental Materials [55].
 In Eq.
-(8), the expectation values are taken over the aforementioned stationary state, while in Eq.
+(8), the expectation values are taken over the stationary state, while in Eq.
 (9) $|\bar{\psi}^{(1)}\rangle$ and $|\bar{\psi}^{(2)}\rangle$ denote projections of the steady-state wavefunction to the 1- and 2-photon subspaces respectively.
 Using Eq.
-(9), we can construct the following account of how photon antibunching arises.
+(9), we construct the following account of how photon antibunching arises.
 By fine-tuning the system parameters, it may be possible to satisfy $\langle i|\bar{\psi}^{(1)}\rangle=0$, which we call a 1-photon dark state, for some cavity $i$.
 In the absence of nonlinearity, such a condition also implies a 2-photon dark state $\langle i,j|\bar{\psi}^{(2)}\rangle=0$, since $|\bar{\psi}^{(2)}\rangle=|\bar{\psi}^{(1)}\rangle\otimes|\bar{\psi}^{(1)}\rangle/\sqrt{2}$.
-We then find that even though the numerator and denominator of (9) both vanish, $f_{ij}(z)$ does not vanish; in fact, $g_{ij}^{(2)}=1$ for all $z$.
+We find that even though the numerator and denominator of (9) both vanish, $f_{ij}(z)$ does not vanish; in fact, $g_{ij}^{(2)}=1$ for all $z$.
 A weak nonlinearity, however, can induce a small misalignment between the 1-photon and 2-photon dark states.
@@
 This is our desired photon antibunching point.
+Figure 1: (a) Schematic of zeros (circles) and pole (cross) of $f_{11}(z)$ in the complex $z$ plane, for a two-cavity model.
+UPB occurs at the zeros.
+(b) Schematic of a four-cavity system exhibiting long-lived photon blockade (LLPB).
+(c) Schematic of a photonic structure that can implement the model in (b), using two coupled microring resonators with clockwise and counterclockwise modes of each ring.
+(d) Distribution of zeros (circles) and pole (cross) of $f_{22}(z)$ for the four-cavity model of (b).
+This illustration is not drawn to scale.
+(e) Equal-time two-photon correlation function $g^{(2)}(0)$ for the four-cavity system versus cavity detuning $\Delta$ and loss rate $\gamma$.
+Upper plot: analytical results obtained from Eqs.
+(20)–(22).
+Near the pole (central red spot), there are two zeros of $g^{(2)}(0)$ (blue spots) corresponding to LLPB.
+The model parameters are $k=16$, $J=0.1227$, $J^{\prime}=0.02454$, and $\alpha=0.001227$, chosen so that one of the LLPB points occurs at $\gamma=1$.
+Lower plot: WFMC simulation results, obtained by directly solving Lindblad master equation with the same parameters and $F_{d}=10^{-5}$.
 It is also helpful to introduce the 1-photon and non-interacting 2-photon Green’s functions,
@@
 Eqs.
-(10)–(12) predict antibunching in the driven cavity ($d=1$), $g_{11}^{(2)}(0)=0$, if and only if $z^{3}\approx-\alpha J^{2}/2$ [53].
-The three roots correspond to three 2-photon dark states, and are depicted by blue circles in Fig.
+(10)–(12) predict antibunching in the driven cavity, $g_{11}^{(2)}(0)=0$, if and only if $z^{3}\approx-\alpha J^{2}/2$ [55].
+The three roots, corresponding to three 2-photon dark states, are depicted by blue circles in Fig.
 1(a).
-They lie close to the 1-photon dark state at $z=0$, for which $\langle 1|\bar{\psi}^{(1)}\rangle\propto G_{11}=z(z^{2}-J^{2})^{-1}=0$, as indicated by the red cross in the figure.
+They lie close to the 1-photon dark state at $z=0$, for which $\langle 1|\bar{\psi}^{(1)}\rangle\propto G_{11}=z(z^{2}-J^{2})^{-1}=0$, marked by a red cross in the figure.
 We focus on the root with positive cavity loss, for which
@@
 The fast oscillation in $\cos(J\tau)$ yields a small time window of $\delta\tau\sim 1/J$, tied to the large coupling eigenvalues ±$\pm J$.
-Photon antibunching in this model follows the mechanism laid out by Eqs.
-(8)–(12).
-The 2-photon dark state is perturbed away from the 1-photon dark state, and since the latter occurs at $z=0$, the solution we arrive at involves small $\gamma$ (relative to $J$).
+As laid out by Eqs.
+(8)–(12), photon antibunching occurs when the 2-photon dark state is perturbed away from the 1-photon dark state; since the latter occurs at $z=0$, the solution involves small $\gamma$ (relative to $J$).
 The lattice model studied in Ref.
 52 also works similarly.
-Figure 1: (a) Schematic of the zeros (circles) and pole (cross) of $f_{11}(z)$ in the complex $z$ plane, for a two-cavity model.
-The pole occurs at $z=0$, and UPB occurs at the zeros.
-(b) Schematic of a four-cavity setup that can exhibit long-lived photon blockade (LLPB).
-(c) Distribution of zeros (circles) and pole (cross) of $f_{22}(z)$ for the four-cavity model of (b).
-This illustration is not drawn to scale.
-(d) Equal-time two-photon correlation function $g^{(2)}(0)$ for the four-cavity system versus cavity detuning $\Delta$ and loss rate $\gamma$.
-Upper plot: analytical results obtained from Eqs.
-(20)–(22).
-Near the pole (central red spot), there are two zeros of $g^{(2)}(0)$ (blue spots) corresponding to LLPB.
-The model parameters are $k=16$, $J=0.1227$, $J^{\prime}=0.02454$, and $\alpha=0.001227$, chosen so that one of the LLPB points occurs at $\gamma=1$.
-Lower plot: WFMC simulation results, obtained by directly solving Lindblad master equation with the same parameters and $F_{d}=10^{-5}$.
-Long-Lived Photon Blockade—The above scenario can be bypassed if we find a configuration of cavities with a 1-photon dark state at large $\gamma$.
-Specifically, we want the linear single-photon Green’s function to satisfy $G_{id}(z)=0$ for some $i,d$, with $\gamma=-2\mathrm{Im}(z)\gg\mathrm{max}\,|J_{ij}|$.
-This turns out to be impossible for $N=2$ and $N=3$ (though if we relax the assumption that the cavities are identical, we can find a zero for $N=3$ if one cavity is completely lossless); for details, see the Supplemental Materials [53].
-For $N=4$, however, the condition can be satisfied.
-Consider the configuration shown in Fig.
+Long-Lived Photon Blockade—We now present a system with a 1-photon dark state at large $\gamma$.
+We want the linear single-photon Green’s function to satisfy $G_{id}(z)=0$ for some $i,d$, with $\gamma=-2\mathrm{Im}(z)\gg\mathrm{max}\,|J_{ij}|$, but this turns out to be impossible for $N=2$ and $N=3$ (if we relax the assumption that the cavities are identical, there can be a zero for $N=3$ if one cavity is completely lossless); for details, see the Supplemental Materials [55].
+For $N=4$, however, the condition can be satisfied with the setup in Fig.
 1(b), with four cavities arranged in a ring with couplings $J_{12},J_{23},J_{34},J_{41}\in\mathbb{R}^{+}$; the driven cavity is $d=1$, and photons are extracted from cavity $i=2$.
+This configuration can be implemented using photonic structures like the one shown in Fig.
+1(c), with two ring resonators each hosting a pair of weakly-coupled clockwise and counterclockwise modes [56, 57, 58, 59].
+In the Supplemental Materials, we present further details about this exemplary coupled-ring setup, including full-wave simulations showing how it can achieve the desired model parameters [55].
 A Dyson series expansion in the weak-coupling regime [60] indicates that $G_{12}(z)=G_{21}(z)$ can have a zero that meets the above conditions if $J_{12}\ll J_{23},J_{34},J_{41}$.
@@
 $\delta z\approx\pm{\frac{\sqrt{38}}{16}}\sqrt{\alpha\gamma}e^{-i\pi/4}.$ (22)
-LLPB is predicted to occur at $z\approx z_{0}\pm\delta z$, as shown schematically in Fig. 1(c).
-In the upper plot of Fig.
-1(d), we plot $g_{22}^{(2)}(0)$ versus the cavity parameters $\Delta$ and $\gamma$ using Eqs.
+Therefore, LLPB occurs at $z\approx z_{0}+\delta z$, as shown schematically in Fig. 1(d).
+In the upper panel of Fig.
+1(e), we plot $g_{22}^{(2)}(0)$ versus the cavity parameters $\Delta$ and $\gamma$ using Eqs.
 (20)–(22).
 The parameters, stated in the figure caption, satisfy $\alpha\ll\gamma$ (weak nonlinearity) and $J^{\prime}/k\ll J^{\prime}\ll J\ll\gamma$ (weak coupling).
-For comparison, the lower plot of Fig.
-1(d) shows the results obtained based on wavefunction Monte Carlo (WFMC) [55, 56, 57, 58] simulations, which directly solve the Lindblad master equation (4) using $F_{d}=10^{-5}$, with all other parameters kept the same.
-It is worth noting that the WFMC simulations give $g_{22}^{(2)}(0)\ll 1$, but not exactly zero, in the weak-drive regime.
-This value of $g_{22}^{(2)}(0)$ increases with $F_{d}$, alongside an increase in the mean photon occupation number in the signal cavity [53].
+The lower panel of Fig.
+1(e) shows results from wavefunction Monte Carlo (WFMC) [61, 62, 63, 64] simulations that directly solve the Lindblad master equation (4), using $F_{d}=10^{-5}$ with all other parameters kept the same.
+Note that the WFMC simulations give $g_{22}^{(2)}(0)\ll 1$, but not exactly zero, in the weak-drive regime; the value of $g_{22}^{(2)}(0)$ increases with $F_{d}$, alongside an increase in the mean photon occupation number in the signal cavity [55].
 Figure 2: Time-delayed second-order correlation function $g^{(2)}(\tau)$ for a four-cavity system with LLPB (green curve), a two-cavity system with UPB (blue curve), and a one-cavity system with conventional photon blockade (orange dashes).
@@
 It is remarkable that shortening the cavity lifetime can effectively lengthen the antibunching time window, which can be regarded as a fresh example of non-Hermiticity (in the form of dissipation) producing unexpected physical behaviors [65, 66].
-Our proposal does not need one-way inter-cavity couplings or other exotic requirements [51, 39], and should be highly compatible with quantum optics experiments.
+Our proposal does not need one-way inter-cavity couplings or other exotic requirements [53, 41], and should be compatible with quantum optics experiments using setups like the one shown in Fig.
+1(c) [55].
 Our theoretical framework offers ample scope for further generalization and optimization.
@@
 (b) Signal cavity photon occupation number versus equal-time second order correlation.
+S6 Photonics Realization
+In this section, we explore how the quantum optical model could be implemented in practice.
+To achieve the four coupled optical modes depicted in Fig.
+1(b) of the main text, we can employ either four coupled microring resonators, as shown in Fig.
+S5(a), or two coupled microring resonators with clockwise (CW) and counter-clockwise (CCW) modes, as shown in Fig.
+S5(b) and Fig.
+1(c) of the main text.
+We will focus on the latter case.
+Figure S5: Photonic implementations of the model.
+(a) Configuration using four coupled microring resonators.
+(b) Configuration using two coupled microring resonators where the clockwise (CW) and counter-clockwise (CCW) modes within each ring are coupled.
+(c) Inter-ring coupling strength $J$ as a function of the gap size $d_{gap}$ between the two rings, calculated using Finite Element Method (FEM) simulations.
+In the two-ring setup, the CW and CCW modes have the same natural frequencies.
+They are coupled to each other, with coupling rates $J^{\prime}$ and $J^{\prime\prime}$, which can be achieved by deliberately-introduced subwavelength defects or surface roughness, as shown in previous studies [56, 59, 58, 57].
+Assuming the two rings have the same shape and size, the Langevin equation for the system is given by
+$i\frac{d}{dt}\begin{pmatrix}a_{1}\\ a_{2}\\ a_{3}\\ a_{4}\end{pmatrix}=\begin{pmatrix}\Delta-\frac{i}{2}(\gamma+\gamma_{in}+\gamma_{out})&J^{\prime\prime}&0&J\\ J^{\prime\prime}&\Delta-\frac{i}{2}(\gamma+\gamma_{in}+\gamma_{out})&J&0\\ 0&J&\Delta-\frac{i}{2}\gamma&J^{\prime}\\ J&0&J^{\prime}&\Delta-\frac{i}{2}\gamma\end{pmatrix}\begin{pmatrix}a_{1}\\ a_{2}\\ a_{3}\\ a_{4}\end{pmatrix}+\begin{pmatrix}i\sqrt{\gamma_{\textrm{in}}}\alpha_{\textrm{in}}\\ 0\\ 0\\ 0\end{pmatrix},$ (S59)
+where $\alpha_{\textrm{in}}$ is the input field amplitude, $\gamma$ is the intrinsic decay rate of the rings, $\gamma_{\textrm{in}}$ is the input coupling rate, and $\gamma_{\textrm{out}}$ is the output coupling rate.
+The output field is given by $\alpha_{\textrm{out}}=\sqrt{\gamma_{\textrm{out}}}a_{2}$.
+The parameters $J$, $J^{\prime}$, and $J^{\prime\prime}$ represent the coupling strength between the two rings, the coupling strength between the CW and CCW modes within the second ring, and the coupling strength between the CW and CCW modes within the first ring, respectively.
+The detuning $\Delta=\omega_{c}-\omega_{L}$ is the difference between the cavity resonance frequency $\omega_{c}$ and the laser frequency $\omega_{L}$.
+(We set $\hbar=1$.)
+The Hamiltonian terms describing the self-Kerr and cross-Kerr nonlinearities (between CW and CCW modes) are
+$\mathcal{H}_{nl}=\sum_{i}\alpha a_{i}^{\dagger}a_{i}^{\dagger}a_{i}a_{i}+2\alpha a_{1}^{\dagger}a_{1}a_{2}^{\dagger}a_{2}+2\alpha a_{3}^{\dagger}a_{3}a_{4}^{\dagger}a_{4},$ (S60)
+where $\alpha$ is the Kerr coefficient of each ring resonator.
+We perform Finite Element Method (FEM) simulations of this setup, using exemplary parameters drawn from the experimental literature.
+We assume each ring is composed of silicon carbide (SiC), with radius $R=3\,\mu$m, waveguide thickness $t=0.35$ $\mu$m, and waveguide width $w=0.8\,\mu$m, operating at a wavelength near $\lambda=1550$ nm.
+The refractive indices are $n_{\textrm{SiC}}=2.45$ for the waveguides and $n(\text{SiO}_{2})=1.44$ for the cladding.
+The nonlinear refractive index for SiC is $n_{2}=4.8\times 10^{-6}$ $\mu$m2/W.
+The nonlinearity parameter $\alpha$ is estimated by [11, 38]
+$\alpha=\frac{c\hbar^{2}\omega_{c}^{2}}{n_{\textrm{SiC}}^{2}V}n_{2},$ (S61)
+where $V$ is the mode volume, approximated here by the volume of the microring resonator $V=2\pi Rwt$. This yields $\alpha\approx 4.7\times 10^{-6}$ $\mu$eV.
+Figure S5(c) shows the coupling strength $J$ as a function of the gap size $d_{\textrm{gap}}$ between the two rings, derived from FEM simulations.
+In these simulations, the eigenenergy splitting between the symmetric and antisymmetric supermodes of the coupled system corresponds to $2J$, which is the eigenvalue difference of the matrix
+$\begin{pmatrix}\Delta&J\\ J&\Delta\end{pmatrix}.$
+The FEM simulation identifies an eigenmode at $\lambda=1553$ nm with a simulated intrinsic quality factor $Q_{\textrm{int}}\approx 1.5\times 10^{6}$.
+Accounting for other potential loss sources, we estimate the total intrinsic loss rate of the cavity as $\gamma=5$ $\mu$eV, corresponding to an unloaded quality factor $Q_{\textrm{cav}}=\hbar\omega_{c}/\gamma\approx 1.6\times 10^{5}$.
+The input and output coupling rates are chosen to be $\gamma_{\textrm{in}}=\gamma_{\textrm{out}}=2.5$ $\mu$eV.
+The couplings strengths $J^{\prime}$ and $J^{\prime\prime}$, can be controlled either by adjusting the gap size $d_{\textrm{gap}}$ in the four-cavity configuration [Fig.
+S5(a)] or by introducing a subwavelegnth defect or slightly increase surface roughness [56, 58, 57, 59], that weakly couple the CW and CCW modes in each ring for the two-cavity setup (Fig.
+S5(b)).
+To achieve LLPB, we require $J^{\prime\prime}\ll J^{\prime}$, but their actual values do not need to be precisely controlled; thus, we take $J^{\prime}=0.2$ $\mu$eV and $J^{\prime\prime}=0.02$ $\mu$eV.
+Since experimental parameters may deviate from these nominal values, the detuning $\Delta=\hbar(\omega_{c}-\omega_{L})$ and the inter-ring coupling $J$ are treated as tunable parameters.
+The detuning $\Delta$ can be adjusted by tuning the laser frequency $\omega_{L}$.
+Fine-tuning the coupling $J$ can be achieved by first selecting a gap size $d_{\textrm{gap}}$ corresponding to the desired approximate value of $J$ (from Fig.
+S5(c)) and then continuously adjusting it.
+This fine-tuning can be accomplished by illuminating the gap region with a control laser source.
+Modulating the intensity of this control source allows for precise regional adjustments of the refractive index.
+Figure S6(b) illustrates the dependence of $J$ on the regional refractive index change $\Delta n$ within a spot of radius $r=0.5$ $\mu$m located in the gap (schematically shown in Fig.
+S6(a)), assuming an initial gap size of $d_{\textrm{gap}}=1.74$ $\mu$m.
+Figure S6(c) displays a colormap of the zero-delay second-order correlation function $g^{(2)}(0)$ as a function of the detuning $\Delta$ and coupling strength $J$.
+The black contour line indicates the boundary of the antibunching region, defined by $g^{(2)}(0)<0.5$.
+The inset shows the time-delayed second-order correlation function $g^{(2)}(\tau)$ for the output field, calculated at the parameter values marked by the purple cross in the main panel.
+It shows a over $1$ ns time window for $g^{(2)}(\tau)<0.5$.
+Figure S6: Fine-tuning and simulation results.
+(a) Schematic illustrating the $r=0.5$ $\mu$m spot within the inter-ring gap used for refractive index-based fine-tuning of the coupling $J$.
+(b) Dependence of the coupling strength $J$ on the regional refractive index change $\Delta n$ applied to the spot shown in (a), assuming a fixed gap size $d_{\textrm{gap}}=1.74$ $\mu$m.
+(c) Colormap of the zero-delay second-order correlation function $g^{(2)}(0)$ for the output field as a function of detuning $\Delta$ and coupling strength $J$.
+The black contour line ($g^{(2)}(0)=0.5$) delineates the antibunching region.
+The inset displays the time-delayed second-order correlation $g^{(2)}(\tau)$ calculated at the parameter point indicated by the purple cross.

### v3

@@
 We present a method to achieve photon blockade over a large time window of several cavity lifetimes, even exceeding that of conventional photon blockade, while still requiring only weak nonlinearity.
-This “long-lived photon blockade” (LLPB) occurs when the single-photon Green’s function exhibits a zero at a large cavity loss rate, which is satisfied by an exemplary configuration of four coupled cavities under weak driving.
+This “long-lived photon blockade” (LLPB) occurs when the single-particle Green’s function exhibits a zero at a large cavity loss rate, which is satisfied by an exemplary configuration of four coupled cavities under weak driving.
 Our analytical results agree well with wavefunction Monte Carlo simulations.
@@
 Introduction—Photon blockade is a quantum effect whereby one photon in a nonlinear resonator blocks the entry of other photons, giving rise to antibunched photon statistics [1, 2, 3, 4].
-Conventionally, photon blockade requires the optical nonlinearity within a cavity to exceed its loss: i.e., $\alpha\gg\gamma$, where $\alpha$ characterizes the strength of the nonlinearity and $\gamma$ is the cavity decay rate.
-This can be interpreted in terms of a two-photon state being driven off-resonance by the nonlinearity, shifting its energy by more than the cavity linewidth [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24].
-Achieving this in the optical regime (including telecom wavelengths) is challenging, as material nonlinearities tend to be very weak relative to cavity linewidths.
-Systems featuring exceptionally strong optical nonlinearities are required, such as exciton-polaritons [25, 26], trions [27], Rydberg excitons [28], and Moire excitons [29].
+Conventionally, photon blockade requires the optical nonlinearity in a cavity to exceed its loss: i.e., $\alpha\gg\gamma$, where $\alpha$ characterizes the strength of the nonlinearity and $\gamma$ is the cavity decay rate.
+This can be interpreted as a two-photon state being driven off-resonance by the nonlinearity, shifting its energy by more than the cavity linewidth [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24].
+Achieving this in the optical regime (including telecom wavelengths) is challenging as optical nonlinearities tend to be weak relative to cavity linewidths.
+Systems with exceptionally strong nonlinearities are required, such as exciton-polaritons [25, 26], trions [27], Rydberg excitons [28], and Moire excitons [29].
 The strong-nonlinearity condition may be bypassed via the unconventional photon blockade (UPB) phenomenon, which uses destructive interference between two or more nonlinear cavities (or cavity modes) to cancel the two-photon amplitudes [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52].
 It is commonly believed that UPB requires the inter-cavity coupling rate $J$ to satisfy
-$J\gg\gamma,$ (1)
-which seems to make intuitive sense: for interference to play a role, the photons must be able to hop many times before leaking away [30].
-In this regime, it then turns out that even a weak nonlinearity, $\alpha\ll\gamma$, can produce photon antibunching.
-However, even if the second-order photon correlation $g^{(2)}(\tau)$ vanishes at $\tau=0$, the suppression only holds over a time window of
+$J\gg\gamma.$ (1)
+The intuition is that for interference to play a role, the photons should be able to hop many times before leaking away [30].
+In this regime, it turns out that even a weak nonlinearity, $\alpha\ll\gamma$, can cause photon antibunching, i.e., the vanishing of the second-order photon correlation $g^{(2)}(\tau)$ at $\tau=0$.
+However, the suppression only holds over a time window
 $\delta\tau\sim 1/J\ll 1/\gamma,$ (2)
-due to the Rabi-like oscillations that $J$ induces in the time-dependent state amplitudes.
-This is much smaller than the antibunching time window of $\delta\tau\sim 1/\gamma$ for conventional photon blockade, and poses a serious practical obstacle to the realization and exploitation of UPB.
-For example, a proposal for UPB at telecom wavelengths [38] requires $J\approx 20\gamma$ and yields an antibunching time window of $\sim 100\,\text{ps}$, around the resolution limit for current photodetectors.
-In such two-cavity models, the UPB condition reduces to $\gamma^{3}\propto\alpha J^{2}$, implying that the cavity quality factors satisfy $Q^{3}\propto\omega_{0}^{3}\delta\tau^{2}/\alpha$, where $\omega_{0}$ is the cavity resonance frequency [32].
-Given a lower bound on $\delta\tau$ imposed by detector resolution limits, this condition sets a minimum requirement on $Q$ that is unfavorable in the optical regime, where $\omega_{0}$ is large and optical nonlinearities are typically weak.
-Flayac and Savona have suggested that with one-way dissipative inter-cavity couplings, the rapid oscillations in $g^{(2)}(\tau)$ can be avoided [53]; however, such couplings are highly challenging to implement experimentally.
-In this paper, we show that interference-assisted photon blockade does not require the strong-coupling condition (1), even with weakly nonlinear cavities and ordinary couplings.
-Using a Green’s function formalism, we analyze the steady-state behavior of coupled nonlinear cavities in the weak-driving regime [32], and show that antibunching solutions appear in the vicinity of 1-photon dark states, where the 1-photon Green’s function vanishes in a specific cavity.
-In the two-cavity systems where UPB was first discovered [30, 31, 32], as well as more complicated coupled-cavity configurations [52], such dark states occur only at $\gamma=0$, giving rise to UPB solutions with nonzero but small $\gamma$ satisfying Eq.
+due to the Rabi-like state amplitude oscillations induced by $J$.
+This is much smaller than the antibunching time window of $\sim 1/\gamma$ for conventional photon blockade, and poses a serious practical obstacle to the realization of UPB.
+For example, a proposal for UPB at telecom wavelengths [38] requires $J\approx 20\gamma$ and yields a time window of $\delta\tau\sim 100\,\text{ps}$, which is near the resolution limit of current photodetectors.
+In two-cavity systems, the UPB condition can be reduced to $\gamma^{3}\propto\alpha J^{2}$, and the required cavity quality factors are $Q^{3}\propto\omega_{0}^{3}\delta\tau^{2}/\alpha$, where $\omega_{0}$ is the cavity resonance frequency [32].
+With a lower bound on $\delta\tau$ imposed by detector resolution limits, this sets a minimum requirement on $Q$ that is unfavorable in the optical regime, where $\omega_{0}$ is large and optical nonlinearities are typically weak.
+Flayac and Savona have suggested that one-way dissipative inter-cavity couplings can suppress the rapid oscillations in $g^{(2)}(\tau)$ [53], but such couplings are very challenging to implement experimentally.
+In this paper, we show that interference-assisted photon blockade does not inherently require the strong-coupling condition (1), even if we use ordinary couplings.
+In systems of weakly-nonlinear and weakly-driven cavities [32], photon blockade emerges perturbatively from linear limit solutions of the system that we call “single-particle dark state” (SPDS), defined by the vanishing of the single-particle Green’s function in a designated signal cavity.
+In the two-cavity systems where UPB was first discovered [30, 31, 32], as well as other cavity configurations [52], it so happens that the SPDS only exists under zero loss ($\gamma=0$), so the antibunching solutions generated perturbatively from them satisfy Eq.
 (1).
-However, certain coupled-cavity configurations host other families of dark states at large $\gamma$ (relative to $J$).
-We demonstrate this with a four-cavity system, showing that it produces a qualitatively different form of antibunching that we call long-lived photon blockade (LLPB).
-Unlike the earlier proposal involving one-way dissipative couplings [53], our four-cavity setup involves only standard (i.e., two-way and Hermitian) inter-cavity couplings.
-The antibunching condition $g^{(2)}(0)=0$ is satisfied alongside weak coupling ($J\ll\gamma$) and weak nonlinearity ($\alpha\ll\gamma$).
-Moreover, $g^{(2)}(\tau)$ is suppressed over a time window $\delta\tau\approx 8/\gamma$, two orders of magnitude wider than a comparable two-cavity system exhibiting UPB, and even larger than for conventional photon blockade under strong nonlinearity.
-Thus, it may be possible to achieve photon blockade in the visible regime with weakly nonlinear optical media.
-Model—We consider coupled nonlinear cavities, indexed by $i=1,\dots$, under a weak coherent drive.
+However, the intuition behind Eq.
+(1) is not reliable, as there is no general principle that forbids destructive interference from occurring in lossy systems.
+As an example, we design a system of four lossy cavities ($\gamma\gg J$) hosting an SPDS, via the exact cancellation of multiple strongly-attenuated interference pathways.
+In the weakly-nonlinear quantum regime ($\alpha\ll\gamma$), this gives rise to a qualitatively different form of antibunching we call long-lived photon blockade (LLPB).
+The two-photon correlation function $g^{(2)}(\tau)$ is suppressed over a time window $\delta\tau\approx 8/\gamma$, two orders of magnitude wider than a comparable two-cavity system exhibiting UPB, and even larger than for conventional photon blockade under strong nonlinearity.
+Notably, our setup uses only couplings of the standard (reciprocal and Hermitian) type.
+Using these ideas, it may be possible to achieve photon blockade using photonic devices with weak optical nonlinearities.
+Model—Consider a general system of coupled nonlinear cavities, indexed by $i=1,\dots$, under a weak coherent drive.
 Assuming for simplicity that all the cavities have identical parameters, the cavity Hamiltonian in the frame co-rotating with the driving field, excluding losses, is
@@
 Here, $\hbar=1$, $\Delta$ is the cavity detuning relative to the driving frequency, $\alpha$ is the nonlinearity strength (Kerr coefficient), $J_{ij}$ is the coupling between cavities $i$ and $j$, and $a_{i}$ is the photon annihilation operator for cavity $i$.
-We assume no magneto-optic activity, so $J_{ij}=J_{ji}\in\mathbb{R}$ with a suitable gauge.
-We assume a single-input, single-output setup where a cavity $d$ is driven, and we look at the photon populations in a “readout” cavity (which may or may not be $d$).
-The driving Hamiltonian is $\mathcal{H}_{d}=F_{d}a_{d}^{\dagger}+F^{*}_{d}a_{d}$, where $F_{d}$ is the complex driving amplitude.
-(The following derivation can also be straightforwardly generalized to the case of multiple coherently driven cavities.) The density matrix $\rho$ is governed by the Lindblad master equation [54]
+There is no magneto-optic activity, so $J_{ij}=J_{ji}\in\mathbb{R}$ in a suitable gauge.
+We assume a single-input, single-output setup; a single cavity $d$ is driven, with driving Hamiltonian $\mathcal{H}_{d}=F_{d}a_{d}^{\dagger}+F^{*}_{d}a_{d}$ where $F_{d}$ is the complex driving amplitude.
+(The following derivation can be straightforwardly generalized to multiple coherently driven cavities.) The density matrix $\rho$ is governed by the Lindblad equation [54]
 $i\frac{d\rho}{dt}=[\mathcal{H}_{0}+\mathcal{H}_{d},\rho]+\frac{i\gamma}{2}\sum_{i}(2a_{i}\rho a^{\dagger}_{i}-a^{\dagger}_{i}a_{i}\rho-\rho a^{\dagger}_{i}a_{i}),$ (4)
@@
 In the weak drive limit [32], stochastic quantum jumps—the first term in parentheses—can be ignored, and the problem reduces to a dissipative Schrödinger equation,
-$i\frac{d|\psi\rangle}{dt}$ $=\big{[}\mathcal{H}(z)+\mathcal{H}_{d}\big{]}|\psi\rangle,$ (5)
-$\mathcal{H}(z)$ $\equiv\sum_{i\neq j}J_{ij}a^{\dagger}_{i}a_{j}+\sum_{i}za^{\dagger}_{i}a_{i}+\alpha\sum_{i}a^{\dagger}_{i}a^{\dagger}_{i}a_{i}a_{i},$ (6)
+$i\frac{d|\psi\rangle}{dt}$ $=\big{[}\mathcal{H}+\mathcal{H}_{d}\big{]}|\psi\rangle,$ (5)
+$\mathcal{H}$ $\equiv\sum_{i\neq j}J_{ij}a^{\dagger}_{i}a_{j}+\sum_{i}za^{\dagger}_{i}a_{i}+\alpha\sum_{i}a^{\dagger}_{i}a^{\dagger}_{i}a_{i}a_{i},$ (6)
 $z$ $\equiv\Delta-i\gamma/2,$ (7)
-where the non-Hermitian effective Hamiltonian $\mathcal{H}(z)$ is derived from (3) by adding loss to each site.
-We seek a solution that is stationary (i.e., co-rotating with the drive, $d/dt\rightarrow 0$).
-Truncating to superpositions of only 1- and 2-photon states yields the following equal-time second-order correlation function:
-$g_{ij}^{(2)}$ $=\frac{\langle a^{\dagger}_{j}a^{\dagger}_{i}a_{i}a_{j}\rangle}{\langle a^{\dagger}_{j}a_{j}\rangle\langle a^{\dagger}_{i}a_{i}\rangle}=\big{|}f_{ij}(z)\big{|}^{2},$ (8)
-$f_{ij}(z)$ $=\frac{\sqrt{2}\langle i,j|\bar{\psi}^{(2)}\rangle}{\langle i|\bar{\psi}^{(1)}\rangle\langle j|\bar{\psi}^{(1)}\rangle},$ (9)
-where $|i\rangle=a_{i}^{\dagger}|\text{vac}\rangle$ denotes the state with one photon in cavity $i$ and $|i,j\rangle\equiv|i\rangle\otimes|j\rangle$.
-The full derivation, which generalizes a previous result based on a specific lattice model [52], is given in the Supplemental Materials [55].
+where the non-Hermitian effective Hamiltonian $\mathcal{H}$ is derived from (3) by adding loss to each site.
+The second-order correlation function between sites $i$ and $j$, with time delay $\tau\geq 0$, is defined as
+$g^{(2)}_{ij}(\tau)=\lim\limits_{t\rightarrow+\infty}\frac{\langle a^{\dagger}_{j}(t)a^{\dagger}_{i}(t+\tau)a_{i}(t+\tau)a_{j}(t)\rangle}{\langle a^{\dagger}_{j}(t)a_{j}(t)\rangle\langle a^{\dagger}_{i}(t+\tau)a_{i}(t+\tau)\rangle},$ (8)
+where the expectation values are taken over the vacuum state.
+For photon blockade, we are only concerned with self-correlation ($i=j$), but we leave $i,j$ arbitrary for now.
+To derive (8), we develop the perturbative steady-state solution for (5) to first order in $\alpha$.
+The derivation is presented in the Supplemental Materials [55], and uses the eigendecomposition of the coupling matrix $[J_{ij}]|\varphi_{n}\rangle=\epsilon_{n}|\varphi_{n}\rangle$, where each $|\varphi_{n}\rangle$ is a single-particle eigenmode and $\epsilon_{n}$ is the corresponding eigenenergy relative to the driving frequency.
+We also define the two-photon basis $|\varphi_{mn}\rangle=|\varphi_{m}\rangle\otimes|\varphi_{n}\rangle$.
+The result is
+$g^{(2)}_{ij}(\tau)$ $=\left|1-e^{-iz\tau}\,\frac{2\alpha}{G_{id}G_{jd}}\sum_{k}G_{kd}^{2}G^{(2)}_{ijkk}(\tau)\right|^{2},$ (9)
+$G_{ij}$ $\equiv\sum_{n}\frac{\langle i|\varphi_{n}\rangle\langle\varphi_{n}|j\rangle}{z+\epsilon_{n}},$ (10)
+$G^{(2)}_{ijkk}(\tau)$ $\equiv\sum_{mn}\frac{\langle i,j|\varphi_{mn}\rangle\langle\varphi_{mn}|k,k\rangle}{2z+\epsilon_{m}+\epsilon_{n}}e^{-i\epsilon_{n}\tau}.$ (11)
+Here, $|i\rangle\equiv a_{i}^{\dagger}|\varnothing\rangle$ is the state with one photon in cavity $i$; $|i,j\rangle\equiv|i\rangle\otimes|j\rangle$; and $G\equiv\sum_{n}|\varphi_{n}\rangle\langle\varphi_{n}|/(z+\epsilon_{n})$ is the single-particle Green’s function for the linear system.
+The time-dependent correlation function (9) applies to any configuration of coupled cavities with weak nonlinearity and weak driving, and has not been previously reported to our knowledge.
+From it, we see that the $\tau$-dependence of the correlation has two sources.
+The first is the $e^{-iz\tau}$ factor, which simply accounts for the intrinsic frequency detuning and decay of the cavities; if $\Delta$ is small (i.e., the cavities are driven close to resonance), this reduces to $\approx e^{-\gamma\tau/2}$.
+This $\tau$-dependence is consistent with the conventional photon blockade result $g^{(2)}(\tau)\approx(1-e^{-\gamma\tau/2})^{2}$ [55] (note, however, that conventional blockade uses strong nonlinearity, $\alpha\gg\gamma$).
+The remaining $\tau$-dependence in (9) is from $G^{(2)}_{ijkk}(\tau)$, representing interferometric effects.
 In Eq.
-(8), the expectation values are taken over the stationary state, while in Eq.
-(9) $|\bar{\psi}^{(1)}\rangle$ and $|\bar{\psi}^{(2)}\rangle$ denote projections of the steady-state wavefunction to the 1- and 2-photon subspaces respectively.
-Using Eq.
-(9), we construct the following account of how photon antibunching arises.
-By fine-tuning the system parameters, it may be possible to satisfy $\langle i|\bar{\psi}^{(1)}\rangle=0$, which we call a 1-photon dark state, for some cavity $i$.
-In the absence of nonlinearity, such a condition also implies a 2-photon dark state $\langle i,j|\bar{\psi}^{(2)}\rangle=0$, since $|\bar{\psi}^{(2)}\rangle=|\bar{\psi}^{(1)}\rangle\otimes|\bar{\psi}^{(1)}\rangle/\sqrt{2}$.
-We find that even though the numerator and denominator of (9) both vanish, $f_{ij}(z)$ does not vanish; in fact, $g_{ij}^{(2)}=1$ for all $z$.
-A weak nonlinearity, however, can induce a small misalignment between the 1-photon and 2-photon dark states.
-Given a 1-photon dark state at $z_{0}$, the 2-photon dark state condition is shifted to a nearby value of $z_{0}+\delta z$.
-At this point, the numerator of (9) vanishes while the denominator is nonzero.
-This is our desired photon antibunching point.
+(11), this is written as a superposition of Rabi-like oscillations with eigenfrequencies $\epsilon_{n}$.
+The photon antibunching time-window is therefore determined by $\max\{|\epsilon_{n}|\}$.
 Figure 1: (a) Schematic of zeros (circles) and pole (cross) of $f_{11}(z)$ in the complex $z$ plane, for a two-cavity model.
@@
 Lower plot: WFMC simulation results, obtained by directly solving Lindblad master equation with the same parameters and $F_{d}=10^{-5}$.
-It is also helpful to introduce the 1-photon and non-interacting 2-photon Green’s functions,
-$G(z)$ $\equiv\sum_{n}\frac{|\varphi_{n}\rangle\langle\varphi_{n}|}{z+\epsilon_{n}},$ (10)
-$G^{(2)}(z)$ $\equiv\sum_{mn}\frac{|\varphi_{m},\varphi_{n}\rangle\langle\varphi_{m},\varphi_{n}|}{2z+\epsilon_{m}+\epsilon_{n}},$ (11)
-where $\{|\varphi_{n}\rangle\}$ is the eigenbasis of the coupling matrix, $\left[J_{ij}\right]|\varphi_{n}\rangle=\epsilon_{n}|\varphi_{n}\rangle$, and $|\varphi_{m},\varphi_{n}\rangle\equiv|\varphi_{m}\rangle\otimes|\varphi_{n}\rangle$.
-Using these and assuming $\alpha$ is perturbative, Eq.
-(9) can be expressed as
-$f_{ij}(z)$ $=1-2\alpha\sum_{k}\frac{G_{kd}^{2}}{G_{id}G_{jd}}\langle i,j|G^{(2)}|k,k\rangle.$ (12)
-The time-delayed correlation function $g_{ij}^{(2)}(\tau)$ can be calculated by upgrading the $a_{i}$’s and $a_{i}^{\dagger}$’s to Heisenberg picture operators [55].
-For $\tau>0$, we find that the non-interacting 2-photon Green’s function in Eq.
-(12) is simply replaced by
-$G^{(2)}(z,\tau)$ $\equiv e^{-iz\tau}\sum_{mn}\frac{|\varphi_{m},\varphi_{n}\rangle\langle\varphi_{m},\varphi_{n}|}{2z+\epsilon_{m}+\epsilon_{n}}e^{-i\epsilon_{n}\tau}.$ (13)
-This analytical form of $g_{ij}^{(2)}(\tau)$ has not previously been reported, to our knowledge. Below, we will use it to revisit UPB, and then demonstrate the LLPB phenomenon.
-Conventional blockade—First, let us consider photon blockade in one cavity. In this case, we can bypass the above perturbative procedure and directly find the Green’s functions, obtaining [55]
-$g_{11}^{(2)}(\tau)=\frac{1}{1+4\alpha^{2}/\gamma^{2}}+\frac{4\alpha^{2}/\gamma^{2}}{1+4\alpha^{2}/\gamma^{2}}\left(1-e^{-\frac{\gamma}{2}\tau}\right)^{2}.$ (14)
-The condition $g_{11}^{(2)}(0)\approx 0$ requires $\alpha\gg\gamma$.
-The corresponding time window, estimated as the interval over which $g^{(2)}(\tau)<0.5$, is $\delta\tau=-4\rm{ln}(1-\sqrt{0.5})/\gamma\approx 5/\gamma$.
-UPB with two cavities—For a two-cavity system, with $J_{12}=J_{21}=J\in\mathbb{R}^{+}$, the coupling matrix has eigenvalues $\epsilon_{1,2}=\pm J$ and the eigenbasis $|\varphi_{1,2}\rangle=2^{-1/2}(|1\rangle\pm|2\rangle)$.
-Eqs.
-(10)–(12) predict antibunching in the driven cavity, $g_{11}^{(2)}(0)=0$, if and only if $z^{3}\approx-\alpha J^{2}/2$ [55].
-The three roots, corresponding to three 2-photon dark states, are depicted by blue circles in Fig.
-1(a).
-They lie close to the 1-photon dark state at $z=0$, for which $\langle 1|\bar{\psi}^{(1)}\rangle\propto G_{11}=z(z^{2}-J^{2})^{-1}=0$, marked by a red cross in the figure.
+UPB with two cavities—To illustrate the use of this formalism, consider the two-cavity UPB model with $J_{12}=J_{21}=J\in\mathbb{R}^{+}$, for which the eigenmodes are $|\varphi_{1,2}\rangle=(|1\rangle\pm|2\rangle)/\sqrt{2}$, with $\epsilon_{1,2}=\pm J$.
+Plugging these into Eqs.
+(9)–(11), we find that the antibunching condition $g_{11}^{(2)}(0)=0$ is satisfied if and only if $z^{3}\approx-\alpha J^{2}/2$ [55].
+In Fig.
+1(a), the three roots for $z$ are plotted as blue circles.
+They are symmetrically distributed around the origin (marked by a red cross), the significance of which will be explained shortly.
 We focus on the root with positive cavity loss, for which
 $\gamma=\sqrt{3}\Delta\propto J\sqrt{\alpha/\gamma}.$ (12)
-Unlike the one-cavity case, this is achievable for $\alpha\ll\gamma$, but only if we enter the strong-coupling regime (1), as previously noted [30, 32]. Moreover, from Eq. (13),
-$g_{11}^{(2)}(\tau)$ $\approx\left(1-e^{-\frac{\gamma}{2}\tau}\cos(\Delta\tau)\cos(J\tau)\right)^{2}.$ (16)
-The fast oscillation in $\cos(J\tau)$ yields a small time window of $\delta\tau\sim 1/J$, tied to the large coupling eigenvalues ±$\pm J$.
-As laid out by Eqs.
-(8)–(12), photon antibunching occurs when the 2-photon dark state is perturbed away from the 1-photon dark state; since the latter occurs at $z=0$, the solution involves small $\gamma$ (relative to $J$).
-The lattice model studied in Ref.
-52 also works similarly.
-Long-Lived Photon Blockade—We now present a system with a 1-photon dark state at large $\gamma$.
-We want the linear single-photon Green’s function to satisfy $G_{id}(z)=0$ for some $i,d$, with $\gamma=-2\mathrm{Im}(z)\gg\mathrm{max}\,|J_{ij}|$, but this turns out to be impossible for $N=2$ and $N=3$ (if we relax the assumption that the cavities are identical, there can be a zero for $N=3$ if one cavity is completely lossless); for details, see the Supplemental Materials [55].
-For $N=4$, however, the condition can be satisfied with the setup in Fig.
-1(b), with four cavities arranged in a ring with couplings $J_{12},J_{23},J_{34},J_{41}\in\mathbb{R}^{+}$; the driven cavity is $d=1$, and photons are extracted from cavity $i=2$.
-This configuration can be implemented using photonic structures like the one shown in Fig.
-1(c), with two ring resonators each hosting a pair of weakly-coupled clockwise and counterclockwise modes [56, 57, 58, 59].
-In the Supplemental Materials, we present further details about this exemplary coupled-ring setup, including full-wave simulations showing how it can achieve the desired model parameters [55].
-A Dyson series expansion in the weak-coupling regime [60] indicates that $G_{12}(z)=G_{21}(z)$ can have a zero that meets the above conditions if $J_{12}\ll J_{23},J_{34},J_{41}$.
-Specifically, we consider
-$[J_{ij}]=\begin{pmatrix}0&J^{\prime}/k&0&J\\ J^{\prime}/k&0&J&0\\ 0&J&0&J^{\prime}\\ J&0&J^{\prime}&0\end{pmatrix},$ (17)
-where $k\gg 1$. Using Eqs. (6) and (10),
-$G_{21}(z)$ $=-\frac{J^{\prime}}{k}(z+z_{0})(z-z_{0})/\det\mathcal{H},$ (18)
-$z_{0}$ $\equiv-iJ\sqrt{k-(J^{\prime}/J)^{2}}.$ (19)
-At $z=z_{0}$, $G_{21}$ has a zero and hence $f_{22}$ has a pole (we ignore $z=-z_{0}$, which involves putting gain in the cavities).
-From Eqs.
-(7) and (19), we have $\Delta=0$ and $\gamma\approx 2\sqrt{k}J$, consistent with the weak-coupling regime.
-Next, we search near $z_{0}$ for a zero of $f_{22}(z)$. Substituting Eq. (18) back into Eq. (12), we find that for $z\approx z_{0}$,
-$f(z)$ $\approx\frac{(z-z_{0}-\delta z)(z-z_{0}+\delta z)}{(z-z_{0})^{2}},$ (20)
-$\delta z^{2}$ $=-\dfrac{k^{2}\alpha}{2z_{0}^{2}J^{\prime 2}}\left[{\sum}_{i}G^{(2)}_{22ii}\,\big{(}G_{i1}\,\cdot\,\det\mathcal{H}\big{)}^{2}\right]_{z=z_{0}}$ (21)
-where $G^{(2)}_{22ii}\equiv\langle 2,2|G^{(2)}(z)|i,i\rangle$. By further assuming $J\gg J^{\prime}$, it is approximated as [55]
-$\delta z\approx\pm{\frac{\sqrt{38}}{16}}\sqrt{\alpha\gamma}e^{-i\pi/4}.$ (22)
-Therefore, LLPB occurs at $z\approx z_{0}+\delta z$, as shown schematically in Fig. 1(d).
+Unlike conventional photon blockade, this is achievable for $\alpha\ll\gamma$, but only in the strong-coupling regime (1) [30, 32]. Moreover, from Eqs. (9)–(11) we can derive
+$g_{11}^{(2)}(\tau)$ $\approx\left(1-e^{-\frac{\gamma}{2}\tau}\cos(\Delta\tau)\cos(J\tau)\right)^{2},$ (13)
+yielding a short time window of $\delta\tau\sim 1/J$.
+Single-particle dark state—The condition for photon antibunching in a designated signal cavity $s$ is $g^{(2)}_{ss}(0)=0$. Using Eq. (9), this can be re-expressed as
+$f_{ss}(z)\equiv\frac{G_{sd}^{2}-2\alpha\sum_{k}G_{kd}^{2}G^{(2)}_{sskk}(0)}{G_{sd}^{2}}=0.$ (14)
+The Green’s functions, and hence $f_{ss}(z)$, are meromorphic functions in the complex $z$ plane.
+We claim that for small $\alpha$, the zeros of $f_{ss}(z)$ lie close to the zeros of $G_{sd}(z)$.
+To see why, observe that $f_{ss}(z)=1$ over the whole $z$-plane when $\alpha=0$.
+For $\alpha>0$, $f_{ss}$ develops poles at discrete points $\{z_{n}\}$ which are the zeros of $G_{sd}(z)$.
+By Cauchy’s argument principle, the net number of zeros and poles (counted with their multiplicities and orders) near each $z_{n}$ is conserved, so one or more zeros of $f_{ss}(z)$ must emerge continuously from $z_{n}$.
+This completes the argument.
+The zeros of $G_{sd}(z)$, which we call SPDSs, are instances where the linear version of the system experiences complete destructive interference in the signal cavity.
+Motivated by the above discussion, we seek a system with an SPDS in the $\gamma\gg\max\{|\epsilon_{n}|\}$ regime.
+Since $\max\{|\epsilon_{n}|\}$ is the spectral radius of $[J_{ij}]$, which for an $N\times N$ matrix satisfies $\max\{J_{ij}\}\leq\max\{\absolutevalue{\epsilon_{n}}\}\leq N\max\{J_{ij}\}$, we can reformulate the criterion as
+[…]
+Long-Lived Photon Blockade—Fig.
+1(b) shows an exemplary four-cavity setup that can satisfy (15).
+The cavities are arranged in a ring with couplings $J_{12},J_{23},J_{34},J_{41}\in\mathbb{R}^{+}$.
+The driven cavity is $d=1$, and the signal cavity is $s=2$.
+To understand intuitively how the SPDS arises, we use a Dyson series in the variable $1/z$, which is the Green’s function for an isolated cavity.
+Each term in the series can be interpreted as a photon trajectory starting from $d$ and ending at $s$ [56]; since we intend $\gamma$ to be large relative to the coupling rates, long trajectories that hop multiple times are negligible.
+The Dyson series for $G_{21}$ to order $1/z^{4}$ contains the following five trajectories:
+Trajectory Dyson series term
+$1\mapsto 2$ $\frac{1}{z}\cdot J_{12}\cdot\frac{1}{z}$
+$1\mapsto 2\mapsto 1\mapsto 2$ $\frac{1}{z}\cdot J_{12}\cdot\frac{1}{z}\cdot J_{12}\cdot\frac{1}{z}\cdot J_{12}\cdot\frac{1}{z}$
+$1\mapsto 2\mapsto 3\mapsto 2$ $\frac{1}{z}\cdot J_{12}\cdot\frac{1}{z}\cdot J_{23}\cdot\frac{1}{z}\cdot J_{23}\cdot\frac{1}{z}$
+$1\mapsto 4\mapsto 1\mapsto 2$ $\frac{1}{z}\cdot J_{41}\cdot\frac{1}{z}\cdot J_{41}\cdot\frac{1}{z}\cdot J_{12}\cdot\frac{1}{z}$
+$1\mapsto 4\mapsto 3\mapsto 2$ $\frac{1}{z}\cdot J_{41}\cdot\frac{1}{z}\cdot J_{34}\cdot\frac{1}{z}\cdot J_{23}\cdot\frac{1}{z}$
+Equating the sum of these terms to zero, we obtain [55]
+$z=\pm i\sqrt{J_{41}^{2}+J_{12}^{2}+J_{23}^{2}+\frac{J_{41}J_{23}J_{34}}{J_{12}}}.$ (16)
+By choosing $J_{12}\ll J_{14},J_{23},J_{34}$, we can get a root with $\gamma=-2\mathrm{Im}(z)\approx 2(J_{41}J_{23}J_{34}/J_{12})^{1/2}$, satisfying Eq. (15).
+This four-cavity setup appears to be a minimal configuration for achieving LLPB.
+Applying the above Dyson series analysis (up to second-lowest order in $1/z$) to two- and three-cavity systems, we find no configurations supporting an SPDS at large $\gamma$.
+For two cavities, the Dyson series yields $z=\pm iJ_{12}$, whereas for three cavities it gives $z=-J_{12}-J_{13}$ or $z=-J_{13}J_{23}/J_{12}$; none of these solutions satisfy the criterion (15) [55].
+We now adopt a mirror-symmetric configuration $J_{12}=J^{\prime}/k$, $J_{14}=J_{23}=J$, and $J_{34}=J^{\prime}$.
+The above condition for the SPDS can then be met by choosing $J\gtrsim J^{\prime}$ and $k\gg 1$.
+One potential advantage is that this can be realized with the structure shown in Fig.
+1(c), consisting of two ring resonators each hosting a pair of weakly-coupled clockwise and counterclockwise modes [57, 58, 59, 60].
+In the Supplemental Materials, we present further details about this coupled-ring setup, including full-wave simulations showing how to achieve the desired model parameters [55].
+If we further take $J\gg J^{\prime}$, we can obtain the closed-form approximation
+$f_{22}(z)$ $\approx\frac{(z-z_{0}-\delta z)(z-z_{0}+\delta z)}{(z-z_{0})^{2}},$ (17)
+$z_{0}\approx$ $-i\sqrt{k}J,\hskip 10.0pt\delta z\approx\pm{\frac{\sqrt{38}}{16}}\sqrt{\alpha\gamma}e^{-i\pi/4}.$ (18)
+The SPDS occurs at $z=z_{0}$, with $|\mathrm{Im}(z_{0})|=\sqrt{k}J\gg J$, thus satisfying Eq.
+(15).
+This point, along with the nearby zeros of $f_{22}(z)$—where LLPB is predicted to occur—are depicted in Fig.
+1(d).
 In the upper panel of Fig.
-1(e), we plot $g_{22}^{(2)}(0)$ versus the cavity parameters $\Delta$ and $\gamma$ using Eqs.
-(20)–(22).
-The parameters, stated in the figure caption, satisfy $\alpha\ll\gamma$ (weak nonlinearity) and $J^{\prime}/k\ll J^{\prime}\ll J\ll\gamma$ (weak coupling).
-The lower panel of Fig.
-1(e) shows results from wavefunction Monte Carlo (WFMC) [61, 62, 63, 64] simulations that directly solve the Lindblad master equation (4), using $F_{d}=10^{-5}$ with all other parameters kept the same.
-Note that the WFMC simulations give $g_{22}^{(2)}(0)\ll 1$, but not exactly zero, in the weak-drive regime; the value of $g_{22}^{(2)}(0)$ increases with $F_{d}$, alongside an increase in the mean photon occupation number in the signal cavity [55].
+1(e), we plot $g_{22}^{(2)}(0)$ against $\Delta$ and $\gamma$, calculated from the crude approximation Eqs.
+(17)–(18) using the system parameters stated in the figure caption.
+Note that the vertical axis is flipped relative to Fig.
+1(d).
+We observe strong local minima, consistent with the above predictions.
+For comparison, the lower panel of Fig.
+1(e) shows results from wavefunction Monte Carlo (WFMC) [61, 62, 63, 64] simulations that directly solve the Lindblad equation (4) for the same cavity parameters, using $F_{d}=10^{-5}$ and Fock cutoffs chosen per site to balance accuracy and efficiency.
+Details about the simulations may be found in the Supplemental Materials [55].
+One notable difference is that the WFMC simulations have $g_{22}^{(2)}(0)\ll 1$ but not exactly zero; this residual value increases with $F_{d}$, alongside an increase in the mean photon occupation number in the signal cavity [55].
 Figure 2: Time-delayed second-order correlation function $g^{(2)}(\tau)$ for a four-cavity system with LLPB (green curve), a two-cavity system with UPB (blue curve), and a one-cavity system with conventional photon blockade (orange dashes).
@@
 Antibunching time window—In Fig.
-2, the green curve shows the time-delayed second-order correlation function $g^{(2)}(\tau)$ for the above four-cavity system, at one of its LLPB points (at $\gamma=1$).
-The photon antibunching time window is approximately $8/\gamma$.
-This plot is obtained from WFMC simulations, but the analytical theory gives almost identical results.
-The analytic approximation for this result can be derived for $J^{\prime}\ll J$ by approximating the eigenvalues and eigenvectors of the coupling Hamiltonian to second order in $J^{\prime}$.
-The full result is given in Supplemental Materials [55].
-If we look into the short time scale $\tau\sim 1/\gamma$ and expand $g^{(2)}_{22}(\tau)$ as polynomial of $\gamma\tau$, assuming negligible $\alpha$, we find
-$g_{22}^{(2)}(\tau)$ $\approx(\gamma\tau)^{4}/64.$ (23)
-This small $\tau$ expansion is order-4 and fundamentally distinct from the quadratic conventional photon blockade case $g^{(2)}(\tau)\approx(\gamma\tau)^{2}/4$, by Eq. (14).
+2, we plot the time-delayed correlation function $g^{(2)}(\tau)$ versus $\tau$, for the above four-cavity system at the $\gamma=1$ LLPB point (green curve).
+The photon antibunching time window, which we characterize by $g^{(2)}(\tau)<0.5$, is $\delta\tau\approx 8/\gamma$.
+This plot is obtained from WFMC simulations, but the theory based on Eqs.
+(9)–(11) gives almost identical results [55].
 For comparison, the blue curve in Fig.
-2 shows $g^{(2)}(\tau)$ for a comparable two-cavity UPB setup [30].
-Here we take $\alpha=0.001227$, the same as in the four-cavity LLPB case.
-For the same decay rate $\gamma=1$, two-cavity UPB is achieved for $J\approx 17.67$, and the strong coupling results in rapid oscillations of $g^{(2)}(\tau)$.
-The central antibunching window, estimated as the window over which $g^{(2)}(\tau)<0.5$, is around 90 times smaller than in the LLPB case.
+2 shows $g^{(2)}(\tau)$ for an analogous two-cavity UPB setup [30].
+With the same nonlinearity strength $\alpha$ and decay rate $\gamma$ used in the four-cavity LLPB result, the two-cavity system achieves UPB for $J\approx 17.67$.
+This results in rapid oscillations of $g^{(2)}(\tau)$, with an antibunching window (estimated in the same way) around 90 times smaller than under LLPB.
 As a further comparison, the orange dashes in Fig.
 2 show $g^{(2)}(\tau)$ for a one-cavity setup with $\gamma=1$ and $\alpha=10$.
-(Note that $g^{(2)}(0)$ does not reach precisely zero under conventional photon blockade; we simply choose $\alpha$ so that it is visibly suppressed.) The $g^{(2)}(\tau)$ curves for both LLPB and conventional photon blockade have no oscillations, unlike the two-cavity UPB case.
-Remarkably, the LLPB curve has a significantly flatter bottom near $\tau=0$ than the curve for conventional photon blockade, and the time window is larger by a factor of around 1.68.
-This can be understood from Eq.
-(23); the expansion leading term is of order-4 in $\gamma\tau$, approximately $(\gamma\tau)^{4}/64$, compared to the quadratic term $\approx(\gamma\tau)^{2}/4$ for the single-cavity case.
-This characteristic results in a more gradual change in the dip, which may be advantageous for pulsed device operations [38, 41].
+(Note that $g^{(2)}(0)$ does not reach precisely zero under conventional photon blockade; we simply choose $\alpha$ so that it is visibly suppressed.)
+Returning to the LLPB result, we also note that $g^{(2)}_{22}(\tau)$ scales very slowly for small $\tau$.
+To quantify this, we expand it as a polynomial in $\gamma\tau$ for $\gamma\tau\ll 1$.
+We obtain
+$g_{22}^{(2)}(\tau)$ $\approx(\gamma\tau)^{4}/64,$ (19)
+which is indeed much slower than the UPB and conventional photon blockade results, both of which scale as $\tau^{2}$ [55].
+This “flat bottom” feature of LLPB may be particularly advantageous for pulsed device operations [38, 41].
 Conclusion—We have re-analyzed the origins of unconventional photon blockade (UPB), in which photon antibunching is achieved by quantum interference in weakly nonlinear coupled cavities [30].
 A major problem with UPB has hitherto been the rapid oscillation in $g^{(2)}(\tau)$ due to strong inter-cavity coupling [38, 41].
-We show that this can be overcome by designing a coupled-cavity system so that the linear single-photon Green’s function exhibits a zero (a 1-photon dark state) at a large cavity loss rate.
+We show that this can be overcome by designing a coupled-cavity system so that the linear single-particle Green’s function exhibits a zero (an SPDS) at a large cavity loss rate.
 This leads to a “long-lived photon blockade” (LLPB) scenario whereby $g^{(2)}(\tau)\propto\tau^{4}$ and the time window exceeds that of conventional photon blockade under strong nonlinearity.
-It is remarkable that shortening the cavity lifetime can effectively lengthen the antibunching time window, which can be regarded as a fresh example of non-Hermiticity (in the form of dissipation) producing unexpected physical behaviors [65, 66].
+It is remarkable that shortening the cavity lifetime can effectively lengthen the antibunching time window, which can be regarded as a fresh example of non-Hermiticity (in the form of dissipation) producing unexpected physical effects [65, 66].
 Our proposal does not need one-way inter-cavity couplings or other exotic requirements [53, 41], and should be compatible with quantum optics experiments using setups like the one shown in Fig.
 1(c) [55].
-Our theoretical framework offers ample scope for further generalization and optimization.
+This theoretical framework offers ample scope for further generalization and optimization.
 For instance, in the four-cavity configuration of Fig.
@@
 $=\frac{1}{\sqrt{2}}|\bar{\psi}^{(1)}\rangle\otimes|\bar{\psi}^{(1)}\rangle-\sqrt{2}\alpha F_{d}^{2}\sum_{k}G_{kd}^{2}G^{(2)}|k,k\rangle,$ (S6)
-where $G_{kd}\equiv\langle k|G|d\rangle$ is the $k,d$ element of the single-photon Green’s function and
+where $G_{kd}\equiv\langle k|G|d\rangle$ is the $k,d$ element of the single-particle Green’s function and
 $G^{(2)}\equiv\sum_{mn}\frac{|\varphi_{nm}\rangle\langle\varphi_{mn}|}{2z+\epsilon_{m}+\epsilon_{n}}$ (S7)
@@
 Finally, plugging Eq. (S17) and Eq. (S16) back into Eq. (S15) gives
-$\langle i|\psi_{j}^{(1)}(\tau)\rangle=\langle j|\bar{\psi}^{(1)}\rangle\langle i|\bar{\psi}^{(1)}\rangle-2\alpha F_{d}^{2}\sum_{k}G_{kd}^{2}\langle i,j|G^{(2)}(\tau)|k,k\rangle,\$ (S18)
+$\langle i|\psi_{j}^{(1)}(\tau)\rangle=\langle j|\bar{\psi}^{(1)}\rangle\langle i|\bar{\psi}^{(1)}\rangle-2\alpha F_{d}^{2}e^{-iz\tau}\sum_{k}G_{kd}^{2}\langle i,j|G^{(2)}(\tau)|k,k\rangle,\$ (S18)
 where
-$G^{(2)}(\tau)\equiv\sum_{mn}\frac{|\varphi_{mn}\rangle\langle\varphi_{mn}|}{2z+\epsilon_{m}+\epsilon_{n}}e^{-i(z+\epsilon_{n})\tau}.$ (S19)
+$G^{(2)}(\tau)\equiv\sum_{mn}\frac{|\varphi_{mn}\rangle\langle\varphi_{mn}|}{2z+\epsilon_{m}+\epsilon_{n}}e^{-i\epsilon_{n}\tau}.$ (S19)
 We finally arrive at a perturbative solution for the delayed second-order correlation function:
 $g_{ij}^{(2)}(\tau)$ $=\left|\frac{\langle i|\psi_{j}^{(1)}(\tau)\rangle^{2}}{\bar{n}_{i}\bar{n}_{j}}\right|=\left|\frac{\langle i|\psi_{j}^{(1)}(\tau)\rangle}{\langle i|\bar{\psi}^{(1)}\rangle\langle j|\bar{\psi}^{(1)}\rangle}\right|^{2}$ (S20)
-$=\left|1-2\alpha\sum_{k}\frac{G_{kd}^{2}}{G_{id}G_{jd}}\langle i,j|G^{(2)}(\tau)|k,k\rangle\right|^{2}.$ (S21)
+$=\left|1-2\alpha e^{-iz\tau}\sum_{k}\frac{G_{kd}^{2}}{G_{id}G_{jd}}\langle i,j|G^{(2)}(\tau)|k,k\rangle\right|^{2}.$ (S21)
 S2 Conventional photon blockade
@@
 First, we will develop some intuition about how photon antibunching arises in this system.
-In the Dyson series expansion for $G_{id}(z)$, each term in the series can be interpreted as a photon trajectory starting at the driven cavity $d$ and ending at the output cavity $i$ [60].
+In the Dyson series expansion for $G_{id}$ as a function of $z$, each term in the series can be interpreted as a photon trajectory starting at the driven cavity $d$ and ending at the output cavity $i$ [56].
 Along a trajectory, each cavity contributes $1/z$ to the Green’s function.
@@
 (b) Signal cavity photon occupation number versus equal-time second order correlation.
-S6 Photonics Realization
+S6 WFMC simulation details
+Wave function Monte Carlo (WFMC) simulations [61, 62, 63, 64] were performed using the QuantumOptics.jl toolbox [67], operating in the tensor product of truncated Fock spaces across lattice sites.
+Due to weak driving and uneven photon occupations, we adopted the following strategies to ensure accuracy and efficiency:
+  • To avoid undersampling rare but potentially important jump events, we modified the standard master equation. For jump operators $L_{j}=\sqrt{\gamma}a_{j}$, we applied the transformation:
+    $\mathcal{H}$ $\rightarrow\mathcal{H}^{\prime}=\mathcal{H}+\frac{\beta}{2i}\sum_{j}(L_{j}-L_{j}^{\dagger})$ (S59)
+    $L_{j}$ $\rightarrow L^{\prime}_{j}=L_{j}+\beta\mathcal{I}$ (S60)
+    with $\beta=0.1$. This increases jump rates while preserving the correct ensemble evolution and is known as choosing a different unraveling [54].
+  • Simulations were performed with tolerances $\texttt{RelTol}=\texttt{AbsTol}=10^{-21}$.
+  • To capture steady-state behavior, each trajectory was evolved for a relaxation time $T_{\text{relax}}$, followed by a recording period $T_{\text{record}}$.
+We used $T_{\text{relax}}=100$, $T_{\text{record}}=1000$, and the number of trajectories $N_{\text{traj}}=10$.
+  • Fock cutoffs $N_{\text{Fock}}^{(j)}$ were chosen per site to balance accuracy and efficiency.
+For the driven cavity, we used $N_{\text{Fock}}^{(1)}=16$, while for the other cavities, we used $N_{\text{Fock}}^{(j\neq 1)}=8$.
+Convergence was verified in all cases.
+S7 Photonics Realization
 In this section, we explore how the quantum optical model could be implemented in practice.